A Complete Bibliography of the Journal of the Royal Statistical Society, Series A family: 1880–1889

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

09 November 2023 Version 1.02

Title word cross-reference

&c [393, 432, 159, 177].

15th [89, 90, 260]. 16th [37]. 17th [247, 246]. 1870 [149, 273]. 1879 [27]. 1880 [27]. 1882 [131]. 1883 [162]. 1884 [196, 206, 198]. 1885 [236, 246]. 1886 [284]. 1887 [338, 336]. 1889 [432, 434]. 18th [206]. 19th [433, 432]. 1st [298].

20th [163, 394, 393, 162]. 21st [128, 127]. 23rd [236, 236]. 24th [196, 196]. 26th [156, 382]. 27th [120, 423]. 28th [74, 327]. 29th [284, 284].

30th [27]. 31st [74, 120, 156, 196, 236, 423, 327, 382, 284, 27].

6th [337, 336].

9 [423, 327, 382, 284].

Aberdeen [248]. Abroad [306, 208]. Abuse [345, 364]. Accident [75]. Accomodation [299]. Acquisitions [78]. Act [149]. Additions [14, 26, 35, 50, 13, 25, 34, 49, 62, 73, 87, 99, 61, 86, 111, 119, 125, 134, 110, 145, 155, 160, 172, 186, 195, 204, 213, 227, 235, 244, 257, 271, 283, 293, 301, 282, 318, 326, 334, 350, 370, 381, 391, 402, 413, 422, 430, 439]. Address [89, 393, 432, 395, 37, 91, 434, 127, 162, 338, 336, 296, 164, 130, 206, 246, 165, 129, 248, 198]. Adelphi [423, 327, 382, 284].

Administration [384]. Advantages [378]. Affected [131, 177]. Affecting [274]. after [122, 149]. Age [92, 262]. Agricultural [47, 40, 94, 118, 132, 117, 168, 209, 234, 264, 346, 136, 306, 137, 277]. Agriculture [389, 18]. Aid [405]. Aims [355]. Air [177]. Alcohol [114]. Alexander [78]. Alleged [415]. also [74]. America [38, 246]. American [367]. Among [329]. Amputations [122]. Analysis [53]. Anniversary [28, 27, 74, 120, 156, 196, 237, 285, 328]. Annual [43, 154, 236, 427, 327, 321, 131, 130, 284, 383, 424, 423, 382]. Antiseptic [122]. Application [434, 66]. Applied [170, 151]. April [298]. Arbere [219]. Arrangement [305]. Arrangements [177]. Ascertaining [169, 250]. Association [131, 130, 165, 395, 91, 434, 338, 296, 164, 129, 248, 198]. August [279, 91, 129, 198]. Australasian [46, 141, 254]. Austria [233]. Average [159]. Awards [400].

Bad [117]. Banking [357]. Bankruptcy [80]. Banks [210]. Barbary [385]. Basel [372]. Basel-Stadt [372]. Bath [395]. be [405]. before [149, 273]. Being [162, 336, 199, 39, 206, 246]. Belgium [233]. Bengal [167]. Berlin [399]. Best [81]. Between [189]. Big [368]. Bills [47]. Birmingham [296]. Births [250]. Board [42, 322]. Body [92]. Borough [92]. Bounties [360]. Brigade [6, 56, 106, 142, 182, 220, 266, 314, 359, 410]. Britain [297, 231, 175]. British [395, 91, 434, 338, 296, 164, 129, 248, 198, 398, 18, 206, 198]. Bunker [420]. Burlington [236].

C.B [89, 37]. C.B. [206, 246]. Caird [89]. Calculating [289]. Called [289]. Canada [279, 190]. Canadian [356]. Canal [330]. Cantons [372]. Casualties [22, 69, 116, 152, 192, 232, 290]. Catalogue [203, 280]. Cattle [347]. Cause [191]. Causes [188]. Cavalry [17]. Celebration [212]. Census [11, 82, 71, 167, 233, 279, 397, 427, 426, 38]. Centenary [417]. Century [185, 162, 261, 389, 218]. Certain [4]. Chadwick [427]. Chancellor [336]. Change [169]. Changes [396, 176, 4]. Chapels [82]. Chart [23]. Charter [304]. Children [16]. China [95, 343, 215]. Churches [82]. Circulation [269]. City [82]. Class [355, 320]. Classes [185, 399, 162, 261]. Co [297, 355]. Co-operative [297]. Co-Operators [355]. Coal [81, 420, 231, 404, 406]. College [27, 74, 120, 156, 196]. Colonial [274]. Colonies [46, 254, 206]. Commerce [341]. Commercial [5, 20, 68, 55, 105, 138, 178, 217, 263, 311, 358, 356]. Commission [148]. Committee [11, 67, 397]. Commodities [312, 191, 287, 361, 408]. Common [127]. Companies [437, 273]. Comparative [208]. Compared [124, 179, 396, 270]. Comparison [175, 189]. Condition [322, 373, 277]. Conditions [295, 2]. Confederated [190]. Conference [211, 226]. Congregations [82]. Congress [310, 344, 435, 131, 130]. Connection [188]. Considered [188]. Consumed [420]. Consumption [19, 329]. Continuation [273, 174]. Contribution [219]. Corn [41]. Cornelius [252]. Corners [31]. Cornish [385]. Cost [295]. Cotton [175]. Council [27, 74, 120, 156, 196, 236, 423, 327, 382, 284]. Countries [124, 269, 324, 409, 179, 295, 75, 270]. Court [65]. Crisis [234]. Crops [41]. Customs [249].

Dairy [347]. Dangers [231]. Death [104, 409, 147, 188]. Death-Rate [147, 188]. Deaths [250, 75]. Decade [102]. Decay [140]. December [27, 336, 74, 120, 156, 196, 236, 260, 337, 423, 327, 382, 284]. Decline [118, 313, 147, 188]. Defects [354]. Delivered [89, 432, 37, 127, 162, 336, 206, 246, 393]. Demography [151, 310, 344]. Department [130, 165]. Depopulation [415]. Depression [117]. Different [295]. Directors [437]. Discoveries [400]. Discovery [66]. Distress [47]. Distribution [362, 157, 306, 174, 405, 54]. Districts [415, 277]. Division [322]. Divisions [137]. Divorced [202]. Dock [299]. does [95]. Dr. [150, 251, 393, 432, 375]. Drainage [177]. Drug [95]. Duration [147]. During [14, 6, 21, 56, 78, 106, 182, 266, 314, 359, 410, 341, 142, 220, 101, 273, 97]. Duties [324, 409].

Earnings [208]. Earth [45]. East [373]. Economic [239, 395, 91, 434, 338, 101, 417, 129, 198, 296, 164, 248]. Economical [12, 24, 33, 48, 60, 85, 98, 292, 300, 317, 325, 333, 349, 369, 380, 390, 401, 412, 421, 429, 438, 39]. Economy [151, 436, 434, 130, 165]. Education [149, 148, 16, 148]. Edwin [427]. Effect [147, 191, 2, 406]. Effects [324, 360]. Efficiency [208]. Egypt [197]. Eight [19]. Eighteen [92]. Eighth [120, 395]. Electing [64]. Electoral [176]. Element [407]. Elementary [149]. Emigration [9, 59, 109, 153, 193, 223, 276, 323, 377, 419, 218]. Empire [379, 198]. Ended [27, 74, 120, 156, 196, 236, 423, 327, 382, 284]. Ending [27, 74, 120, 156, 196, 236, 423, 327, 382, 284]. England [154, 194, 71, 70, 306, 338, 149, 29, 262, 415, 2, 305, 65, 30]. English [7, 58, 108, 144, 184, 222, 268, 315, 366, 411, 76, 384, 354, 159, 147, 188, 4, 189]. Entrance [27, 74, 120, 156, 196]. Epidemic [131]. Equivalents [199]. Errata [353, 352]. Esq [89, 37]. Esq. [127]. Essay [239]. Established [124]. Estates [65]. Europe [234, 246]. European [179]. Examinations [387]. Exchange [352, 180, 386, 199]. Exchequer [336]. Expenditure [114, 405]. Experiences [177]. Export [113]. Exports [398, 425]. Express [159, 189].

F

[395, 91, 434, 338, 296, 164, 129, 248, 198, 375]. F.R.S. [89, 393, 432, 37]. Faculties [181]. Failures [97]. Fall [191]. Family [181]. far [177]. Farming [347]. Farr [150]. Fate [202]. Fifteen [273]. Fifth [424, 423, 248]. Fiftieth [196]. Fifty [237, 236, 285, 328, 383, 424, 423, 327, 395, 91, 434, 338, 121, 382, 341, 296, 164, 284, 129, 248, 198]. Fifty-Eighth [395]. Fifty-Fifth [424, 423, 248]. Fifty-First [237, 236, 91]. Fifty-Fourth [383, 382, 198]. Fifty-Ninth [434]. Fifty-Second [285, 284, 129]. Fifty-Seventh [338]. Fifty-Sixth [296]. Fifty-Third [328, 327, 164]. Financial [5, 20, 27, 68, 55, 74, 105, 120, 138, 156, 178, 196, 217, 236, 263, 311, 358, 423, 327, 382, 284]. Fire [6, 56, 106, 142, 182, 220, 266, 314, 359, 410]. Fires [6, 56, 106, 142, 182, 220, 266, 314, 359, 410]. First [237, 236, 91]. Fluctuating [2]. Fluctuations [175]. Food [157]. Foreign [95, 84, 242, 340, 274, 206]. Formation [101]. Forthcoming [212]. Forty [28, 74, 120, 156, 27]. Forty-Eighth [120]. Forty-Ninth [156]. Forty-Seventh [74]. Forty-Sixth [28, 27]. Fourth [383, 382, 198]. France [140, 253, 233, 139, 207, 240, 417, 428, 270]. French [124, 313]. Friendly [216]. Front [303]. Further [261].

G [336]. Gaol [92]. Gardens [236].

General

[20, 151, 383, 424, 423, 376, 382, 198, 256]. German [8, 57, 107, 143, 183, 221, 267, 365, 418]. Germany [255, 362, 378, 277]. Gold [169].

Government [133, 405]. Gradients [159]. Great [297, 231, 175]. Growth [92]. Guy [251].

Hackney [373]. Half [185, 162, 261]. Hamlets [322]. Harvest [19]. Harvest-Years [19]. Harvests [117]. Health [177]. Height [92]. Held [27, 74, 120, 196, 236, 423, 327, 395, 434, 338, 131, 382, 296, 164, 130, 165, 129, 248, 156, 91, 284, 198]. Her [432, 393]. Hill [76]. History [5, 20, 68, 55, 105, 138, 178, 217, 241, 263, 311, 358, 65]. Holdings [306]. Home [208, 19]. Hon [336]. Honorary [393, 432]. Horses [17]. Hospitals [122, 177]. House [27, 74, 120, 156, 196]. Howard [39, 177]. Huddersfield [165]. Human [92]. Hundred [121]. Hygiene [310, 344].

II [78]. Illegitimacy [70]. Illustrated [92, 246]. Illustrating [398]. Illustration [66]. Immigration [9, 59, 109, 153, 193, 223, 276, 323, 377, 419, 200]. Imperial [309]. Import [324, 113]. Importance [76]. Imports [425, 19, 175]. Improvement [177]. Improvements [348]. Inaugural [393, 37, 127, 162, 336, 206]. Including [190]. Inclusive [19]. Income [321]. Incomes [362, 388, 396, 336]. Increase [299, 336, 30]. Incumbered [65]. Index [280, 398]. India [76, 148]. Indian [230]. Indictable [29]. Industrial [93, 201, 211, 226]. Industry [103, 341]. Influence [122]. Influencing [274]. Information [1, 36, 51, 88, 100, 126, 135, 161, 173, 205, 214, 245, 258, 294, 302, 335, 351, 392, 403, 431]. Inhabitants [322]. Inland [241]. Inmates [177]. Inquiry [201]. Institute [309, 316, 246]. Institutions [216, 177]. Instrument [289]. Insurance [10]. Interest [407]. International [77, 310, 157, 39, 246, 316, 344, 435, 246]. Inventions [400]. Ireland [93, 231]. Irish [65]. Iron [23, 406]. Italian [101]. Italy [233, 101, 291]. IV [130, 165].

J [336]. James [89, 37]. January [273]. Japan [332]. Jevons [123]. John [39, 177]. Joint [437, 273]. Jubilee [212, 243]. June [27, 196, 236, 284, 74, 120, 156, 196, 236, 423, 327, 382, 284]. Jurisdiction [29, 65].

K.C.M.G [206, 246]. King

[27, 74, 120, 156, 196]. Kingdom

[101, 71, 242, 388, 275, 425, 137, 231, 329, 174, 273, 298, 97, 75].

Labour [208]. Labours [39]. Land

[171, 54]. Large [122]. Last

[185, 162, 261, 101, 341, 233]. Laws

[229, 416]. Legislation [239]. Levied [372]. Liability [273]. Libraries [84]. Library [14, 26, 35, 50, 13, 25, 34, 49, 62, 73, 87, 99, 61, 86, 111, 119, 125, 134, 110, 145, 155, 160, 172, 186, 195, 204, 213, 227, 235, 244, 257, 271, 283, 293, 301, 280, 282, 318, 326, 334, 350, 370, 381, 391, 402, 413, 422, 430, 439, 203]. Licenses [405]. Life [10, 170, 147, 339, 259]. Limited [273]. Listerian [122]. Literature [7, 8, 58, 57, 108, 107, 144, 143, 184, 183, 222, 221, 268, 267, 315, 367, 366, 365, 411, 418]. Liverpool [92]. LL.D. [162, 374]. Lloyd [22, 69, 116, 152, 192, 232, 290]. Local [43, 154, 194, 405]. London [6, 27, 42, 56, 81, 74, 106, 120, 142, 156, 182, 196, 220, 236, 266, 314, 359, 410, 423, 327, 373, 121, 382, 238, 284]. Long [159]. Loss [259].

M.P [336]. made [39]. Majesty [393, 432]. Male [339]. Males [92]. Manchester [338]. Manufactures [398]. Manufacturing [137]. Many [95]. Marine [22, 69, 116, 152, 192, 232, 290]. Maritime [299]. Marriages [250]. Marseilles [224]. Material [338]. Mathematical [288, 357]. Mathematics [434]. Matter [303]. Means [278]. Measurement [307]. Measuring [376]. Media [199]. Medical [265]. Meeting [28, 27, 74, 120, 156, 196, 237, 236, 285, 328, 327, 395, 91, 434, 338, 296, 164, 284, 129, 248, 198, 27, 383, 424, 423, 382]. Memorial [239]. Men [297]. Merchant [77]. Meritorious [400]. Messages [96]. Metal [231]. Metals [269, 199, 287]. Method [169, 288, 66, 53]. Methods [64, 250, 376, 305]. Metropolitan [137, 6, 56, 106, 142, 182, 220, 266, 314, 359, 410]. Migration [229, 416, 277]. Migrations [253]. Mines [385]. Mining [231, 406]. Misadventure [75]. Moderate [336]. Monetary [269, 435, 307]. Money [2]. Montreal [198]. Mortality [265, 122, 66, 320, 329, 331, 4, 31]. Movement [21]. Mr. [104, 115, 252]. Murder [225].

National [114, 130, 165, 270, 131]. Negligence [75]. Neumann [375]. Neumann-Spallart [375]. Newcastle [434]. Newcastle-Upon-Tyne [434]. Newmarch [104, 115, 239]. Newspaper [83, 281]. Nineteenth [218]. Ninth [156, 434]. North [279, 305]. North-West [279]. Note [139, 38]. Notes [12, 24, 33, 48, 13, 25, 34, 49, 60, 85, 98, 61, 86, 110, 292, 300, 282, 317, 325, 333, 349, 369, 380, 390, 401, 412, 421, 429, 438, 159, 261, 199, 385, 286]. Nottingham [131, 130]. November [432, 162, 206, 246, 89, 90, 128, 163, 247, 394, 433, 393, 37, 127]. Number [75]. Numbers [398]. Numerical [66].

Observations [92]. Occupation [262]. Occupations [275, 322, 373]. October [165]. Offences [29]. Offices [133]. Official [79, 332, 363]. Old [324]. Opening [89, 432, 246]. operative [297]. Operators [355]. Opinions [177]. Opium [95]. Organisation [133, 355]. Organizations [297]. Oriental [39]. Other [124, 179, 305, 75, 270, 177]. Our [340, 176].

Paper [131, 174, 273]. Parliamentary [137]. Part [279]. Past [3]. Peace [131]. Peasant [219]. People [275, 373]. Periodical [15, 63, 83, 112, 146, 187, 228, 272, 281, 319, 371, 414]. Periods [29]. Personal [428]. Physician [393, 432]. Plague [39]. Political [151, 436, 434, 39]. Poor [384, 16]. Popular [149]. Population [21, 45, 72, 140, 253, 313, 343, 298, 238, 215, 54, 30, 277]. Port [81]. Ports [368, 299]. Position [179, 291]. Postal [52]. Pox [255, 131, 121]. Practice [158]. Precious [269, 199, 287]. Premiums [10]. Present [179, 298, 291, 52]. Presented [27, 74, 120, 156, 196, 236, 423, 327, 382, 284]. Preservation [177]. President [89, 393, 432, 395, 37, 91, 434, 127, 162, 338, 336, 296, 164, 130, 206, 246, 165, 129, 248, 198]. Press [83, 281]. Pressure [384]. Preventible [259]. Price [81, 19, 191, 175]. Prices [312, 389, 274, 376, 396, 287, 361, 408]. Principal [299]. Prisons [177]. Proceedings [28, 27, 90, 67, 74, 128, 163, 237, 247, 260, 285, 337, 328, 383, 394, 424, 433]. Processes [170]. Produce [398, 19]. Production [269, 139, 207, 240, 342, 136]. Products [157, 76]. Profession [265]. Professor [123, 374]. Progress [103, 185, 162, 261, 338, 355, 101, 273, 341, 286]. Progressive [398, 278, 372]. Promotion [131, 130, 165]. Property [321, 428]. Proportion [289]. Proposed [405]. Proprietorship [219]. Prospective [176]. Provinces [190]. Prussia [233, 388]. Public [436, 177].

Quantitation [352, 386]. Quarter [14]. Queen [393, 432]. Question [420, 404, 52]. Quinquennial [29].

R. [162]. Railway

[32, 124, 170, 356, 354, 295]. Railways [378, 230]. Rate [384, 250, 338, 147, 188, 407]. Rates [10, 4].

Rationale [180]. Read [131, 174, 273]. Record [181]. Records [92]. Redistribution [137]. Reduction [52]. Reference [11, 397, 39]. Regime [124]. Regions [76]. Registrar [256]. Registrar-General [256]. Reign [78]. Relating [177]. Relation [216, 66, 174, 262]. Relations [39]. Relative [378, 122, 231]. Remarks [329, 2]. Remote [31]. Remuneration [211, 226]. Report [27, 11, 74, 120, 156, 196, 241, 236, 397, 423, 327, 382, 284]. Representation [137]. Representative [176]. Representatives [64]. Resources [93]. Results [20, 42, 151, 233]. Retrospective [241]. Returns [43, 40, 15, 94, 63, 132, 112, 154, 168, 146, 209, 194, 187, 228, 264, 272, 346, 319, 371, 414]. Revenue [241, 174]. Review [151, 291, 176, 190]. Right [336]. Rising [2]. Robert [127]. Rooms [27, 74, 120, 156, 196, 423, 327, 336, 382, 284]. Royal [432, 336, 304, 308, 423, 327, 393, 382]. Royalties [406]. Runs [159]. Rural [415, 277]. Russia [21, 78, 54]. Russian [379].

Sale [47]. Saving [170]. Savings [210]. Saxony [388]. Scales [305]. School [42, 322]. Schools [177]. Science [239, 395, 91, 434, 338, 296, 164, 129, 248, 198, 131, 130, 165]. Sea [259]. Season [262]. Second [285, 284, 129]. Section [395, 91, 434, 338, 296, 164, 129, 248, 198]. September [395, 434, 338, 131, 296, 164, 130, 248]. Serving [329]. Session [162, 206, 432, 336, 246]. Sessional [27, 74, 120, 156, 196, 236, 423, 327, 382, 284]. Seven [19]. Seventeenth [389]. Seventh [74, 338]. Several [339]. Sex [262]. Shipping [77, 299, 360]. Short [219]. Silk [175]. Similar [216, 65]. since [101]. Sir [427, 206, 246]. Sixth [28, 344, 131, 296, 130, 27, 310]. Size [306]. Sliding [305]. Small [255, 122, 131, 121]. Smokers [95]. Social [131, 39, 130, 165]. Societies [171, 216]. Society [27, 89, 74, 120, 156, 203, 212, 196, 236, 243, 304, 308, 423, 327, 393, 432, 37, 127, 162, 336, 382, 273, 206, 246, 284]. Some [13, 25, 34, 49, 61, 86, 110, 282, 76, 356, 384, 376, 199, 197, 246, 75, 241]. Somerset [27, 74, 120, 156, 196]. Southampton [129]. Southport [164]. Spain [179]. Spallart [375]. special [39]. Speed [159]. Stadt [372]. Standard [307]. Stanley [123]. State [216]. States [131, 246, 72, 225, 200, 38]. Stations [76]. Statistical [12, 24, 33, 48, 60, 85, 98, 89, 133, 203, 233, 292, 300, 304, 316, 317, 325, 333, 349, 308, 369, 380, 390, 401, 412, 421, 429, 438, 423, 327, 393, 432, 37, 92, 127, 162, 336, 53, 382, 291, 330, 246, 190, 27, 67, 74, 120, 156, 196, 236, 206, 246, 284]. Statistics [22, 46, 32, 80, 77, 69, 84, 83, 79, 116, 141, 152, 192, 210, 232, 254, 225, 290, 281, 345, 332, 364, 388, 379, 363, 17, 139, 340, 216, 395, 76, 356, 18, 136, 384, 354, 91, 288, 387, 434, 127, 113, 338, 102, 158, 320, 166, 174, 176, 296, 164, 197, 246, 129, 97, 248, 198, 215, 65]. Stock [352, 437, 386, 273]. Story [330]. Strand [27, 423, 327, 382, 284, 74, 120, 156, 196]. Strikes [3]. Structural [177]. Study [219]. Subject [39]. Subjects [363]. such [122]. Suez [330]. Suggestions [137, 426, 65]. Suicides [262]. Summary [29, 339]. Supplement [256]. Supplies [177]. Supply [95]. Survey [29]. Survivorship [331]. Switzerland [233, 372]. Sympathy [436]. System [122, 176, 385]. Systems [124].

T [393, 432]. Table [289]. Tables [331, 339]. Tariff [52]. Tariffs [249, 274]. Taxation [43, 154, 194, 174, 372]. Taxes [321]. Telegraph [96, 52]. Telegraphy [44]. Telelogues [96]. Tell [340]. Temperature [66]. Ten [32, 42, 44, 256, 3, 216, 18]. Tenth [38]. Tenure [54]. Term [166]. Terrace [27, 423, 327, 382, 284, 74, 120, 156, 196]. Territorial [78]. Territories [279]. Theatre [236]. Their [322, 76, 159, 199, 157, 406]. Them [189]. Theory [357, 158]. Therein [176]. Third [328, 327, 164, 407]. Thirty [92]. Those [189]. Thrift [216]. Throughout [2]. Tonnage [102]. Tower [322]. Trade [23, 103, 242, 348, 340, 274, 76, 224, 2, 130, 165]. Trades [406]. Traffic [295]. Training [16]. Trains [159, 189]. Transferred [405]. translation [313]. Travelling [170]. Treasury [67]. Tribunal [65]. Troops [329]. Truth [66]. Tuesday [89, 37, 127]. Turkish [79]. Twenty [131, 19, 101, 130, 286]. Twenty-Eight [19]. Twenty-Seven [19]. Twenty-Sixth [131, 130]. Two [121]. Tyne [434].

United [71, 72, 242, 225, 388, 275, 425, 137, 231, 329, 174, 273, 200, 298, 38, 246, 97, 75]. Universal [289]. University [236]. Unlimited [273]. Unsuitable [363]. Upon [434, 2, 114, 122, 147]. Uri [372]. Use [113]. Utility [127].

v [436]. Value [425, 76, 169, 428, 307, 2]. Values [352, 386]. Variation [376]. Variations [425, 250, 307]. Various [269, 409]. Vaud [372]. view [176]. Violence [75]. Vital [17, 246]. Volume [1, 36, 51, 88, 100, 126, 135, 161, 173, 205, 214, 245, 258, 294, 302, 335, 351, 392, 403, 431, 425].

W [206]. W. [236]. Wage [305]. Wage-Arrangement [305]. Wages [399, 385]. Wales [71, 70, 149, 29, 262, 30]. Walford [252]. War [131]. Water [177]. Waterways [378]. Wealth [270]. Weight [92]. West [279, 385]. Wheat [342, 230, 19, 175]. William [104, 123, 150, 251]. Willis [336]. Wine [139, 207, 240]. Work [385]. Workhouses [177]. Working [297, 185, 399, 162, 261, 295, 355, 176]. Works [12, 24, 33, 48, 60, 85, 98, 292, 300, 317, 325, 333, 349, 369, 380, 390, 401, 412, 421, 429, 438]. World [103, 347, 368, 342, 2, 31].

X [375].

Year [40, 9, 6, 22, 27, 32, 42, 94, 59, 56, 69, 74, 132, 109, 106, 116, 120, 154, 168, 153, 142, 152, 156, 209, 193, 182, 194, 196, 223, 220, 232, 236, 256, 264, 276, 266, 323, 314, 377, 359, 419, 410, 423, 327, 216, 18, 382, 284]. Years [21, 3, 425, 92, 121, 19, 101, 273, 341, 97, 286, 44]. York [91].

Zealand [286].

References

Anonymous:1880:VIa

[1] Anonymous. Volume information. Journal of the Statistical Society, 43(1):i– viii, March 1880. CODEN ???? ISSN

                                             0959-5341. URL http://www.jstor.
                                             org/stable/2339128.

Patterson:1880:VMR

[2] R. H. Patterson. Is the value of money rising in England and throughout the world? With remarks on the effect of the fluctuating conditions of trade upon the value of money. Journal of the Statistical Society, 43(1):1–34, March 1880. CO-DEN ???? ISSN 0959-5341. URL http:

                                             //www.jstor.org/stable/2339129.

Bevan:1880:SPT

[3] G. Phillips Bevan. The strikes of the past ten years. Journal of the Statistical Society, 43(1):35–64, March 1880. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339130.

Welton:1880:CCE

[4] Thomas A. Welton. On certain changes in the English rates of mortality. Journal of the Statistical Society, 43(1):65– 94, March 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339131.

Anonymous:1880:FCH

[5] Anonymous. Financial and commercial history of 1879. Journal of the Statistical Society, 43(1):95–109, March 1880. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339132.

Anonymous:1880:FLD

[6] Anonymous. The fires of London during the year 1879, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 43(1):109–114, March 1880. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339133.

Anonymous:1880:EL

[7] Anonymous. English literature in 1879. Journal of the Statistical Society, 43(1): 114–116, March 1880. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339134.

Anonymous:1880:GL

[8] Anonymous. German literature of 1878 and 1879. Journal of the Statistical Society, 43(1):116–117, March 1880. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339135.

Anonymous:1880:EIY

[9] Anonymous. Emigration and immigration in the year 1879. Journal of the Statistical Society, 43(1):117–123, March 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339136.

Anonymous:1880:RLI

[10] Anonymous. Rates of life insurance premiums. Journal of the Statistical Society, 43(1):123–134, March 1880. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339137.

Anonymous:1880:RCR

[11] Anonymous. Report of a committee with reference to the census of 1881. Journal of the Statistical Society, 43(1):134– 139, March 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339138.

Anonymous:1880:NESa

[12] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 43(1):139–143, March 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339139.

Anonymous:1880:NSAa

[13] Anonymous. Notes on some of the additions to the library. Journal of the Statistical Society, 43(1):143–147, March 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339140.

Anonymous:1880:ALD

[14] Anonymous. Additions to the library during the quarter. Journal of the Statistical Society, 43(1):147–158, March 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339141.

Anonymous:1880:PR

[15] Anonymous. Periodical returns. Journal of the Statistical Society, 43(1):159– 182, March 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339142.

Mouat:1880:ETC

[16] Frederic J. Mouat. On the education and training of the children of the poor. Journal of the Statistical Society, 43(2): 183–250, June 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338999.

Balfour:1880:VSC

[17] T. Graham Balfour. Vital statistics of Cavalry horses. Journal of the Statistical Society, 43(2):251–274, June 1880. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339000.

Craigie:1880:TYS

[18] Patrick George Craigie. Ten years’ statistics of British agriculture, 1870–79. Journal of the Statistical Society, 43(2): 275–312, June 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339001.

Lawes:1880:HPI

[19] J. B. Lawes and J. H. Gilbert. On the home produce, imports, consumption, and price of wheat, over twenty-eight (or twenty-seven) harvest-years, 1852– 53 to 1879–80 inclusive. Journal of the Statistical Society, 43(2):313–340, June 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339002.

Anonymous:1880:GRC

[20] Anonymous. General results of the commercial and financial history of 1879. Journal of the Statistical Society, 43(2): 341–355, June 1880. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339003.

Anonymous:1880:MPR

[21] Anonymous. The movement of the population in Russia during the years 1867– 70. Journal of the Statistical Society, 43 (2):356–364, June 1880. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2339004.

Anonymous:1880:LSM

[22] Anonymous. Lloyd’s statistics of marine casualties for the year 1879. Journal of the Statistical Society, 43(2):365– 379, June 1880. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339005.

Anonymous:1880:ITC

[23] Anonymous. An iron trade chart. Journal of the Statistical Society, 43(2):380– 381, June 1880. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339006.

Anonymous:1880:NESb

[24] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 43(2):382–388, June 1880. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339007.

Anonymous:1880:NSAb

[25] Anonymous. Notes on some of the additions to the library. Journal of the Statistical Society, 43(2):388–393, June 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339008.

Anonymous:1880:ALa

[26] Anonymous. Additions to the library. Journal of the Statistical Society, 43(2): 393–404, June 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339009.

Anonymous:1880:RCF

[27] Anonymous. Report of the Council for the financial year ended 31st December, 1879, and for the sessional year ending 30th June, 1880, presented at the Forty-Sixth Anniversary Meeting of the Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College Entrance), Strand, London, on the 30th of June, 1880, and of the proceedings at the meeting. Journal of the Statistical Society, 43(3):405–416, September 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339175.

Anonymous:1880:PFS

[28] Anonymous. Proceedings of the fortysixth anniversary meeting. Journal of the Statistical Society, 43(3):417–422, September 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339176.

Levi:1880:SIS

[29] Leone Levi. A survey of indictable and summary jurisdiction offences in England and Wales, from 1857 to 1876, in quinquennial periods, and in 1877 and 1878. Journal of the Statistical Society, 43(3):423–461, September 1880. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339177.

Williams:1880:IPE

[30] R. Price Williams. On the increase of population in England and Wales. Journal of the Statistical Society, 43(3):462– 508, September 1880. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2339178.

Westergaard:1880:MRC

[31] Harald Westergaard. Mortality in remote corners of the world. Journal of the Statistical Society, 43(3):509–520, September 1880. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339179.

Anonymous:1880:TYRa

[32] Anonymous. Ten years’ railway statistics. Journal of the Statistical Society, 43 (3):521–531, September 1880. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339180.

Anonymous:1880:NESc

[33] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 43(3):531–547, September 1880. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339181.

Anonymous:1880:NSAc

[34] Anonymous. Notes on some additions to the library. Journal of the Statistical Society, 43(3):547–548, September 1880. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339182.

Anonymous:1880:ALb

[35] Anonymous. Additions to the library. Journal of the Statistical Society, 43(3): 548–558, September 1880. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339183.

Anonymous:1880:VIb

[36] Anonymous. Volume information. Journal of the Statistical Society, 43(4):721– 743, December 1880. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339024.

Caird:1880:IAJ

[37] James Caird. The inaugural address of James Caird, Esq., C.B., F.R.S., President of the Statistical Society, delivered on Tuesday the 16th of November, 1880. Journal of the Statistical Society, 43(4):559–572, December 1880. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339025.

Mouat:1880:NTC

[38] F. J. Mouat. Note on the tenth census of the United States of America. Journal of the Statistical Society, 43(4):573–604, December 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339026.

Potter:1880:OPS

[39] Henry Percy Potter. “The Oriental Plague in its Social, Economical, Political, and International Relations, special Reference Being made to the ‘Labours of John Howard on the Subject’ ”. Journal of the Statistical Society, 43(4):605–642, December 1880. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339027.

Anonymous:1880:ARY

[40] Anonymous. The agricultural returns for the year 1880. Journal of the Statistical Society, 43(4):643–664, December 1880. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339028.

Anonymous:1880:CC

[41] Anonymous. The corn crops of 1880. Journal of the Statistical Society, 43(4): 664–670, December 1880. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2339029.

Anonymous:1880:TYRb

[42] Anonymous. Ten years’ results of the London school board. Journal of the Statistical Society, 43(4):670–682, December 1880. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339030.

Anonymous:1880:ALT

[43] Anonymous. The annual local taxation returns of 1878–79. Journal of the Statistical Society, 43(4):683–687, December 1880. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339031.

Anonymous:1880:TYT

[44] Anonymous. Ten Years’s telegraphy. Journal of the Statistical Society, 43(4): 687–690, December 1880. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2339032.

Anonymous:1880:PE

[45] Anonymous. The population of the Earth. Journal of the Statistical Society, 43(4):690–697, December 1880. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339033.

Anonymous:1880:SAC

[46] Anonymous. Statistics of Australasian colonies. Journal of the Statistical Society, 43(4):698–699, December 1880. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339034.

Anonymous:1880:ADB

[47] Anonymous. Agricultural distress and bills of sale. Journal of the Statistical Society, 43(4):700–705, December 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339035.

Anonymous:1880:NESd

[48] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 43(4):705–708, December 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339036.

Anonymous:1880:NSAd

[49] Anonymous. Notes on some additions to the library. Journal of the Statistical Society, 43(4):709–711, December 1880. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339037.

Anonymous:1880:ALc

[50] Anonymous. Additions to the library. Journal of the Statistical Society, 43(4): 711–720, December 1880. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339038.

Anonymous:1881:VIa

[51] Anonymous. Volume information. Journal of the Statistical Society, 44(1):i– viii, March 1881. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2339103.

Williams:1881:QRP

[52] R. Price Williams. The question of the reduction of the present postal telegraph tariff. Journal of the Statistical Society, 44(1):1–30, March 1881. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339104.

Hooper:1881:MSA

[53] Wynnard Hooper. The method of statistical analysis. Journal of the Statistical Society, 44(1):31–49, March 1881. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339105.

Vesselovsky:1881:LTD

[54] A. V. [A. Vesselovsky]. Land tenure and the distribution of the population in Russia. Journal of the Statistical Society, 44(1):50–74, March 1881. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339106.

Anonymous:1881:FCH

[55] Anonymous. Financial and commercial history of 1880. Journal of the Statistical Society, 44(1):75–89, March 1881. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339107.

Anonymous:1881:FLD

[56] Anonymous. The fires in London during the year 1880, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 44(1):89–94, March 1881. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339108.

Anonymous:1881:GL

[57] Anonymous. German literature of 1879 and 1880. Journal of the Statistical Society, 44(1):95, March 1881. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339109.

Anonymous:1881:EL

[58] Anonymous. English literature in 1880. Journal of the Statistical Society, 44(1): 95–97, March 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339110.

Anonymous:1881:EIY

[59] Anonymous. Emigration and immigration in the year 1880. Journal of the Statistical Society, 44(1):97–103, March 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339111.

Anonymous:1881:NESa

[60] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 44(1):103–105, March 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339112.

Anonymous:1881:NSAa

[61] Anonymous. Notes on some of the additions to the library. Journal of the Statistical Society, 44(1):105–108, March 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339113.

Anonymous:1881:ALa

[62] Anonymous. Additions to the library. Journal of the Statistical Society, 44(1): 109–116, March 1881. CODEN ????

ISSN 0959-5341. URL http://www.

jstor.org/stable/2339114.

Anonymous:1881:PR

[63] Anonymous. Periodical returns. Journal of the Statistical Society, 44(1):117– 140, March 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339115.

Droop:1881:MER

[64] H. R. Droop. On methods of electing representatives. Journal of the Statistical Society, 44(2):141–202, June 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339223.

Urlin:1881:HSI

[65] R. Denny Urlin. The history and statistics of the Irish incumbered estates court, with suggestions for a tribunal with similar jurisdiction in England. Journal of the Statistical Society, 44(2):203–234, June 1881. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339224.

Guy:1881:TRM

[66] William A. Guy. On temperature and its relation to mortality: An illustration of the application of the numerical method to the discovery of truth. Journal of the Statistical Society, 44(2):235– 268, June 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339225.

Anonymous:1881:PSC

[67] Anonymous. The proceedings of the Statistical Committee of the Treasury. Journal of the Statistical Society, 44(2): 269–367, June 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339226.

Anonymous:1881:CFH

[68] Anonymous. Commercial and financial history of 1880. Journal of the Statistical Society, 44(2):368–381, June 1881. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339227.

Anonymous:1881:LSM

[69] Anonymous. Lloyd’s statistics of marine casualties for the year 1880. Journal of the Statistical Society, 44(2):381– 393, June 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339228.

Anonymous:1881:IEW

[70] Anonymous. Illegitimacy in England and Wales, 1879. Journal of the Statistical Society, 44(2):394–398, June 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339229.

Anonymous:1881:CEW

[71] Anonymous. Census of England and Wales and of the United Kingdom, 1881. Journal of the Statistical Society, 44(2): 398–413, June 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339230.

Anonymous:1881:PUS

[72] Anonymous. The population of the United States. Journal of the Statistical Society, 44(2):414–417, June 1881. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2339231.

Anonymous:1881:ALb

[73] Anonymous. Additions to the library. Journal of the Statistical Society, 44(2): 417–428, June 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339232.

Anonymous:1881:RCF

[74] Anonymous. Report of the Council for the financial year ended 31st December, 1880, and for the sessional year ending 28th June, 1881, presented at the forty-seventh anniversary meeting of the Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College entrance), Strand, London, on the 28th of June, 1881; and also, of the proceedings at the meeting. Journal of the Statistical Society, 44(3):429–443, September 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339058.

Walford:1881:NDA

[75] Cornelius Walford. On the number of deaths from accident, negligence, violence, and misadventure in the United Kingdom and some other countries. Journal of the Statistical Society, 44(3): 444–527, September 1881. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339059.

Clarke:1881:ESH

[76] Hyde Clarke. The English stations in the Hill regions of India: Their value and importance, with some statistics of their products and trade. Journal of the Statistical Society, 44(3):528–573, September 1881. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2339060.

Anonymous:1881:ISM

[77] Anonymous. International statistics of merchant shipping. Journal of the Statistical Society, 44(3):574–584, September 1881. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2339061.

Anonymous:1881:TAR

[78] Anonymous. The territorial acquisitions of Russia during the reign of Alexander II. Journal of the Statistical Society, 44 (3):584–588, September 1881. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2339062.

Anonymous:1881:TOS

[79] Anonymous. Turkish official statistics. Journal of the Statistical Society, 44(3): 588–590, September 1881. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339063.

Anonymous:1881:BS

[80] Anonymous. Bankruptcy statistics. Journal of the Statistical Society, 44(3): 590–596, September 1881. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339064.

Anonymous:1881:PBC

[81] Anonymous. Price of best coal in Port of London, 1832–80. Journal of the Statistical Society, 44(3):596, September 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339065.

Anonymous:1881:CCC

[82] Anonymous. Census of congregations of the city churches and chapels. Journal of the Statistical Society, 44(3):596–601, September 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339066.

Anonymous:1881:SNP

[83] Anonymous. Statistics of the newspaper and periodical Press. Journal of the Statistical Society, 44(3):602–608, September 1881. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2339067.

Anonymous:1881:SFL

[84] Anonymous. Statistics of foreign libraries. Journal of the Statistical Society, 44(3):608–611, September 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339068.

Anonymous:1881:NESb

[85] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 44(3):611–614, September 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339069.

Anonymous:1881:NSAb

[86] Anonymous. Notes on some additions to the library. Journal of the Statistical Society, 44(3):614–618, September 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339070.

Anonymous:1881:ALc

[87] Anonymous. Additions to the library. Journal of the Statistical Society, 44(3): 618–628, September 1881. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339071.

Anonymous:1881:VIb

[88] Anonymous. Volume information. Journal of the Statistical Society, 44(4):769– 790, December 1881. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339184.

Anonymous:1881:OAJ

[89] Anonymous. The opening address of James Caird, Esq., C.B., F.R.S., President of the Statistical Society, delivered on Tuesday, 15th November, 1881. Journal of the Statistical Society, 44(4): 629–643, December 1881. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339185.

Anonymous:1881:PN

[90] Anonymous. Proceedings on the 15th November, 1881. Journal of the Statistical Society, 44(4):644–648, December 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339186.

Duff:1881:APS

[91] M. E. Grant Duff. Address of the President of Section F, “Economic Science and Statistics”, of the British Association, at the fifty-first meeting, held at York, in August, 1881. Journal of the Statistical Society, 44(4):649–659, December 1881. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2339187.

Danson:1881:SOG

[92] J. T. Danson. Statistical observations on the growth of the human body (males) in height and weight, from eighteen to thirty years of age, as illustrated by the records of the Borough Gaol of Liverpool. Journal of the Statistical Society, 44(4):660–674, December 1881. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339188.

Bevan:1881:IRI

[93] G. Phillips Bevan. The industrial resources of Ireland. Journal of the Statistical Society, 44(4):675–721, December 1881. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339189.

Anonymous:1881:ARY

[94] Anonymous. The agricultural returns for the year 1881. Journal of the Statistical Society, 44(4):722–743, December 1881. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2339190.

Anonymous:1881:OCH

[95] Anonymous. Opium in China: How many smokers does the foreign drug supply? Journal of the Statistical Society, 44(4):743–750, December 1881. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2339191.

Batten:1881:TTM

[96] J. W. Batten. Telelogues and telegraph messages. Journal of the Statistical Society, 44(4):750–751, December 1881. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339192.

Seyd:1881:SFU

[97] Richard Seyd. Statistics of failures in the United Kingdom during the years 1880 and 1881. Journal of the Statistical Society, 44(4):751–755, December 1881. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2339193.

Anonymous:1881:NESc

[98] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 44(4):755–757, December 1881. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2339194.

Anonymous:1881:ALd

[99] Anonymous. Additions to the library. Journal of the Statistical Society, 44(4): 757–768, December 1881. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2339195.

Anonymous:1882:VIa

[100] Anonymous. Volume information. Journal of the Statistical Society, 45(1):i– viii, March 1882. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338946.

Levi:1882:EPI

[101] Leone Levi. The economic progress of Italy during the last twenty years, since the formation of the Italian kingdom in 1861. Journal of the Statistical Society, 45(1):1–36, March 1882. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2338947.

Glover:1882:TSD

[102] John Glover. Tonnage statistics of the decade 1870–80. Journal of the Statistical Society, 45(1):37–81, March 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338948.

Anonymous:1882:WPT

[103] Anonymous. The world’s progress in trade and industry. Journal of the Statistical Society, 45(1):82–114, March 1882. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338949.

Anonymous:1882:DMW

[104] Anonymous. The death of Mr. William Newmarch. Journal of the Statistical Society, 45(1):115–121, March 1882. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338950.

Anonymous:1882:FCH

[105] Anonymous. Financial and commercial history of 1881. Journal of the Statistical Society, 45(1):121–132, March 1882. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2338951.

Anonymous:1882:FLD

[106] Anonymous. The fires in London during the year 1881, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 45(1):132–136, March 1882. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2338952.

Anonymous:1882:GL

[107] Anonymous. German literature of 1880 and 1881. Journal of the Statistical Society, 45(1):137, March 1882. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2338953.

Anonymous:1882:EL

[108] Anonymous. English literature in 1881. Journal of the Statistical Society, 45(1): 137–139, March 1882. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2338954.

Anonymous:1882:EIY

[109] Anonymous. Emigration and immigration in the year 1881. Journal of the Statistical Society, 45(1):139–145, March 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2338955.

Anonymous:1882:NSA

[110] Anonymous. Notes on some additions to the library. Journal of the Statistical Society, 45(1):145–146, March 1882. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2338956.

Anonymous:1882:ALa

[111] Anonymous. Additions to the library.

Journal of the Statistical Society, 45(1): 146–156, March 1882. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2338957.

Anonymous:1882:PR

[112] Anonymous. Periodical returns. Journal of the Statistical Society, 45(1):157– 180, March 1882. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2338958.

Giffen:1882:UIE

[113] R. Giffen. The use of import and export statistics. Journal of the Statistical Society, 45(2):181–296, June 1882. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979272.

Bourne:1882:NEU

[114] Stephen Bourne. The national expenditure upon alcohol. Journal of the Statistical Society, 45(2):297–332, June 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979273.

Anonymous:1882:MN

[115] Anonymous. Mr. Newmarch. Journal of the Statistical Society, 45(2):333, June 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979274.

Anonymous:1882:LSM

[116] Anonymous. Lloyd’s statistics of marine casualties for the year 1881. Journal of the Statistical Society, 45(2):333– 345, June 1882. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979275.

Anonymous:1882:BHA

[117] Anonymous. Bad harvests and agricultural depression. Journal of the Statistical Society, 45(2):346–358, June 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979276.

Anonymous:1882:AD

[118] Anonymous. Agricultural decline. Journal of the Statistical Society, 45(2):359– 371, June 1882. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979277.

Anonymous:1882:ALb

[119] Anonymous. Additions to the library. Journal of the Statistical Society, 45(2): 371–382, June 1882. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979278.

Anonymous:1882:RCF

[120] Anonymous. Report of the Council for the financial year ended 31st December, 1881, and for the sessional year ending 27th June, 1882, presented at the fortyeighth anniversary meeting of the Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College entrance), Strand, London, on the 27th of June, 1882. Journal of the Statistical Society, 45(3):383–398, September 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979318.

Guy:1882:THF

[121] William A. Guy. Two hundred and fifty years of small pox in London. Journal of the Statistical Society, 45(3):399–443, September 1882. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979319.

Burdett:1882:RMA

[122] Henry C. Burdett. The relative mortality, after amputations, of large and small hospitals, and the influence of the antiseptic (Listerian) system upon such mortality. Journal of the Statistical Society, 45(3):444–483, September 1882. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979320.

Anonymous:1882:PWS

[123] Anonymous. Professor William Stanley Jevons. Journal of the Statistical Society, 45(3):484–488, September 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979321.

Anonymous:1882:FRR

[124] Anonymous. The French railway regime compared with the systems established in other countries. Journal of the Statistical Society, 45(3):489–504, September 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979322.

Anonymous:1882:ALc

[125] Anonymous. Additions to the library.

Journal of the Statistical Society, 45(3):

504–518, September 1882. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979323.

Anonymous:1882:VIb

[126] Anonymous. Volume information. Journal of the Statistical Society, 45(4):641– 661, December 1882. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979172.

Giffen:1882:UCS

[127] Robert Giffen. The utility of common statistics. The inaugural address of Robert Giffen, Esq., President of the Statistical Society, delivered on Tuesday, 21st November, 1882. Journal of the Statistical Society, 45(4):519–546, December 1882. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2979173.

Anonymous:1882:PN

[128] Anonymous. Proceedings on the 21st November, 1882. Journal of the Statistical Society, 45(4):547–548, December 1882. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979174.

Sclater-Booth:1882:APS

[129] G. Sclater-Booth. Address of the President of Section F, “Economic Science and Statistics” of the British Association, at the fifty-second meeting, held at Southampton, in August, 1882. Journal of the Statistical Society, 45(4):549–557, December 1882. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2979175.

Price:1882:APD

[130] Bonamy Price. Address of the President of the Department IV, “Economy and Trade” of the National Association for the Promotion of Social Science, at the twenty-sixth annual Congress, held at Nottingham, in September, 1882. Journal of the Statistical Society, 45(4):558– 576, December 1882. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979176.

Guy:1882:SPE

[131] William A. Guy. The small pox epidemic as affected by the states of war and peace. A paper read at the twentysixth annual Congress of the National Association for the Promotion of Social Science, held at Nottingham, in September, 1882. Journal of the Statistical Society, 45(4):577–587, December 1882. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979177.

Anonymous:1882:ARY

[132] Anonymous. The agricultural returns for the year 1882. Journal of the Statistical Society, 45(4):588–605, December 1882. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979178.

Anonymous:1882:OGS

[133] Anonymous. The organisation of government statistical offices. Journal of the Statistical Society, 45(4):606–623, December 1882. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979179.

Anonymous:1882:ALd

[134] Anonymous. Additions to the library.

Journal of the Statistical Society, 45(4): 623–640, December 1882. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979180.

Anonymous:1883:VIa

[135] Anonymous. Volume information. Journal of the Statistical Society, 46(1):i– viii, March 1883. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2979279.

Craigie:1883:SAP

[136] P. G. Craigie. Statistics of agricultural production. Journal of the Statistical Society, 46(1):1–58, March 1883. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979280.

Ellis:1883:PRM

[137] Arthur Ellis. The parliamentary representation of the metropolitan, agricultural, and manufacturing divisions of the United Kingdom, with suggestions for its redistribution. Journal of the Statistical Society, 46(1):59–100, March 1883. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979281.

Anonymous:1883:FCH

[138] Anonymous. Financial and commercial history of 1882. Journal of the Statistical Society, 46(1):101–113, March 1883. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979282.

Bateman:1883:NSW

[139] A. E. Bateman. A note on the statistics of wine production in France. Journal of the Statistical Society, 46(1):113– 119, March 1883. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2979283.

Anonymous:1883:DPF

[140] Anonymous. Decay of population in France. Journal of the Statistical Society, 46(1):119–128, March 1883. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979284.

Anonymous:1883:AS

[141] Anonymous. Australasian statistics.

Journal of the Statistical Society, 46(1): 128–145, March 1883. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979285.

Anonymous:1883:FLD

[142] Anonymous. The fires in London during the year 1882, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 46(1):145–149, March 1883. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979286.

Anonymous:1883:GL

[143] Anonymous. German literature of 1881 and 1882. Journal of the Statistical Society, 46(1):150, March 1883. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979287.

Anonymous:1883:EL

[144] Anonymous. English literature in 1882.

Journal of the Statistical Society, 46(1): 150–152, March 1883. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979288.

Anonymous:1883:ALa

[145] Anonymous. Additions to the library.

Journal of the Statistical Society, 46(1): 152–164, March 1883. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979289.

Anonymous:1883:PR

[146] Anonymous. Periodical returns. Journal of the Statistical Society, 46(1):165– 188, March 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979290.

Humphreys:1883:RDE

[147] Noel A. Humphreys. The recent decline in the English death-rate, and its effect upon the duration of life. Journal of the Statistical Society, 46(2):189– 224, June 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979239.

Johnston:1883:EII

[148] James Johnston. Education in India and the India Commission on Education. Journal of the Statistical Society, 46(2): 225–282, June 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979240.

Hamilton:1883:PEE

[149] Rowland Hamilton. Popular education in England and Wales before and after the Elementary Education Act of 1870. Journal of the Statistical Society, 46(2): 283–349, June 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979241.

Anonymous:1883:DWF

[150] Anonymous. Dr. William Farr. Journal of the Statistical Society, 46(2):350– 351, June 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979242.

Anonymous:1883:RGR

[151] Anonymous. Review of the general results of demography applied to political economy. Journal of the Statistical Society, 46(2):351–367, June 1883. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979243.

Anonymous:1883:LSM

[152] Anonymous. Lloyd’s statistics of marine casualties for the year 1882. Journal of the Statistical Society, 46(2):367– 379, June 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979244.

Anonymous:1883:EIY

[153] Anonymous. Emigration and immigration in the year 1882. Journal of the Statistical Society, 46(2):380–387, June 1883. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979245.

Anonymous:1883:ALT

[154] Anonymous. The annual local taxation returns (england) of the year 1880–81. Journal of the Statistical Society, 46(2): 387–391, June 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979246.

Anonymous:1883:ALb

[155] Anonymous. Additions to the library.

Journal of the Statistical Society, 46(2): 392–404, June 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979247.

Anonymous:1883:RCF

[156] Anonymous. Report of the Council for the financial year ended 31st December, 1882, and for the sessional year ending 26th June, 1883, presented at the fortyninth anniversary meeting of the Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College entrance), Strand, London, on the 26th of June, 1883. Journal of the Statistical Society, 46(3):405–422, September 1883. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979301.

Bourne:1883:FPT

[157] Stephen Bourne. Food products and their international distribution. Journal of the Statistical Society, 46(3):423–460, September 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979302.

Hooper:1883:TPS

[158] Wynnard Hooper. The theory and practice of statistics. Journal of the Statistical Society, 46(3):461–516, September 1883. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979303.

Foxwell:1883:EET

[159] E. Foxwell. English express trains: Their average speed, &c., with notes on gradients, long runs, &c. Journal of the Statistical Society, 46(3):517–574, September 1883. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2979304.

Anonymous:1883:ALc

[160] Anonymous. Additions to the library. Journal of the Statistical Society, 46(3): 574–592, September 1883. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979305.

Anonymous:1883:VIb

[161] Anonymous. Volume information. Journal of the Statistical Society, 46(4):739– 760, December 1883. CODEN ????

ISSN 0959-5341. URL http://www.

jstor.org/stable/2979306.

Giffen:1883:PWC

[162] R. Giffen. The progress of the working classes in the last half century: Being the inaugural address of R. Giffen, LL.D., President of the Statistical Society, session 1883–84. delivered 20th November, 1883. Journal of the Statistical Society, 46(4):593–622, December 1883. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979307.

Anonymous:1883:PN

[163] Anonymous. Proceedings on the 20th November, 1883. Journal of the Statistical Society, 46(4):623–625, December 1883. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979308.

Palgrave:1883:ESS

[164] R. H. Inglis Palgrave. “Economic science and statistics.” The address of the President of Section F of the British Association, at the fifty-third meeting, held at Southport, in September, 1883. Journal of the Statistical Society, 46(4):626–636, December 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979309.

Rogers:1883:ETA

[165] J. E. Thorold Rogers. “Economy and trade.” The address of the President of the Department IV of the National Association for the Promotion of Social Science, held at Huddersfield in October, 1883. Journal of the Statistical Society, 46(4):637–655, December 1883. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979310.

John:1883:TS

[166] V. John. The term “statistics.”. Journal of the Statistical Society, 46(4):656–679, December 1883. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2979311.

Anonymous:1883:CB

[167] Anonymous. Census of Bengal, 1881. Journal of the Statistical Society, 46(4): 680–690, December 1883. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979312.

Anonymous:1883:ARY

[168] Anonymous. The agricultural returns for the year 1883. Journal of the Statistical Society, 46(4):691–713, December 1883. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979313.

Edgeworth:1883:MAC

[169] F. Y. Edgeworth. On the method of ascertaining a change in the value of gold. Journal of the Statistical Society, 46(4):714–718, December 1883. CODEN ???? ISSN 0959-5341. URL http://

                                        www.jstor.org/stable/2979314.

Anonymous:1883:LSP

[170] Anonymous. Life saving processes applied to railway travelling. Journal of the Statistical Society, 46(4):719–720, December 1883. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2979315.

Anonymous:1883:LS

[171] Anonymous. Land societies. Journal of the Statistical Society, 46(4):720–721, December 1883. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979316.

Anonymous:1883:ALd

[172] Anonymous. Additions to the library.

Journal of the Statistical Society, 46(4): 721–738, December 1883. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979317.

Anonymous:1884:VIa

[173] Anonymous. Volume information. Journal of the Statistical Society, 47(1):i– viii, March 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979209.

Levi:1884:SRU

[174] Leone Levi. Statistics of the revenue of the United Kingdom from 1859–82, in relation to the distribution of taxation (in continuation of a paper read in 1860). Journal of the Statistical Society, 47(1):1–33, March 1884. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979210.

Poynting:1884:CFP

[175] J. H. Poynting. A comparison of the fluctuations in the price of wheat and in the cotton and silk imports into Great Britain. Journal of the Statistical Society, 47(1):34–74, March 1884. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979211.

Martin:1884:ESR

[176] John Biddulph Martin. Electoral statistics: A review of the working of our representative system from 1832–1881, in view of prospective changes therein. Journal of the Statistical Society, 47 (1):75–124, March 1884. CODEN ????

ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979212.

Sweeting:1884:EOJ

[177] R. D. R. Sweeting. “The experiences and opinions of John Howard on the preservation and improvement of the health of the inmates of schools, prisons, workhouses, hospitals, and other public institutions, as far as health is affected by structural arrangements relating to supplies of air and water, drainage, &c.”. Journal of the Statistical Society, 47(1): 125–141, March 1884. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979213.

Anonymous:1884:FCH

[178] Anonymous. Financial and commercial history of 1883. Journal of the Statistical Society, 47(1):142–150, March 1884. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979214.

Delboy:1884:PPS

[179] P. A. Delboy. The present position of Spain compared with other European countries. Journal of the Statistical Society, 47(1):150–163, March 1884. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979215.

Edgeworth:1884:RE

[180] F. Y. Edgeworth. The rationale of exchange. Journal of the Statistical Society, 47(1):164–166, March 1884. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979216.

Anonymous:1884:RFF

[181] Anonymous. Record of family faculties. Journal of the Statistical Society, 47(1):

166, March 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979217.

Anonymous:1884:FLD

[182] Anonymous. The fires in London during the year 1883, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 47(1):166–170, March 1884. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979218.

Anonymous:1884:GL

[183] Anonymous. German literature of 1882 and 1883. Journal of the Statistical Society, 47(1):171, March 1884. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979219.

Anonymous:1884:EL

[184] Anonymous. English literature in 1883.

Journal of the Statistical Society, 47(1): 171–173, March 1884. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979220.

Anonymous:1884:PWC

[185] Anonymous. “The progress of the working classes in the last half century. Journal of the Statistical Society, 47(1):174, March 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979221.

Anonymous:1884:ALa

[186] Anonymous. Additions to the library.

Journal of the Statistical Society, 47(1): 175–196, March 1884. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979222.

Anonymous:1884:PR

[187] Anonymous. Periodical returns. Journal of the Statistical Society, 47(1):197– 220, March 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979223.

Longstaff:1884:RDE

[188] G. B. Longstaff. The recent decline in the English death-rate considered in connection with the causes of death. Journal of the Statistical Society, 47(2): 221–258, June 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979188.

Willock:1884:EET

[189] H. B. Willock. English express trains in 1871; and a comparison between them and those of 1883. Journal of the Statistical Society, 47(2):259–311, June 1884. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979189.

Walford:1884:SRC

[190] Cornelius Walford. A statistical review of Canada, including its Confederated Provinces. Journal of the Statistical Society, 47(2):312–347, June 1884. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979190.

Leroy-Beaulieu:1884:FPC

[191] Paul Leroy-Beaulieu. The fall in the price of commodities; its cause and effect. Journal of the Statistical Society, 47(2):348–359, June 1884. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979191.

Anonymous:1884:LSM

[192] Anonymous. Lloyd’s statistics of marine casualties for 1883. Journal of the Statistical Society, 47(2):360–367, June 1884. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979192.

Anonymous:1884:EIY

[193] Anonymous. Emigration and immigration in the year 1883. Journal of the Statistical Society, 47(2):368–374, June 1884. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979193.

Anonymous:1884:LTR

[194] Anonymous. Local taxation returns (england) for the year 1881–82. Journal of the Statistical Society, 47(2):374– 380, June 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979194.

Anonymous:1884:ALb

[195] Anonymous. Additions to the library.

Journal of the Statistical Society, 47(2): 380–394, June 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979195.

Anonymous:1884:RCF

[196] Anonymous. Report of the Council for the financial year ended 31st December, 1883, and for the sessional year ending 24th June, 1884, presented at the fiftieth anniversary meeting of the Statistical Society, held at the Society’s rooms, Somerset House Terrace (King’s College entrance), Strand London, on the 24th of June, 1884. Journal of the Statistical Society, 47(3):395–414, September 1884. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979248.

Rabino:1884:SSE

[197] Joseph Rabino. Some statistics of Egypt. Journal of the Statistical Society, 47(3): 415–467, September 1884. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979249.

Temple:1884:GSB

[198] Richard Temple. “The general statistics of the British empire.” The address of the President of Section F, “Economic Science and Statistics,” of the British Association at the fifty-fourth meeting, held at Montreal, in August, 1884. Journal of the Statistical Society, 47(3):468–484, September 1884. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979250.

Martin:1884:MEB

[199] John Biddulph Martin. Media of exchange: Being some notes on the precious metals and their equivalents. Journal of the Statistical Society, 47(3):485– 495, September 1884. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979251.

Liegeard:1884:IUS

[200] Armand Liegeard. Immigration into the United States. Journal of the Statistical Society, 47(3):496–516, September 1884. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979252.

Harrison:1884:NII

[201] Frederic Harrison. A new industrial inquiry. Journal of the Statistical Society, 47(3):516–519, September 1884. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979253.

Bertillon:1884:FD

[202] Jacques Bertillon. The fate of the divorced. Journal of the Statistical Society, 47(3):519–526, September 1884. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979254.

Anonymous:1884:CLS

[203] Anonymous. Catalogue of the Library of the Statistical Society. Journal of the Statistical Society, 47(3):526–528, September 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979255.

Anonymous:1884:ALc

[204] Anonymous. Additions to the library.

Journal of the Statistical Society, 47(3): 529–546, September 1884. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979256.

Anonymous:1884:VIb

[205] Anonymous. Volume information. Journal of the Statistical Society, 47(4):713– 733, December 1884. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979144.

Rawson:1884:BFC

[206] Rawson W. Rawson. British and Foreign Colonies: Being the inaugural address of Sir Rawson W. Rawson, K.C.M.G., C.B., President of the Statistical Society, session 1884–85. delivered 18th November, 1884. Journal of the Statistical Society, 47(4):547–608, December 1884. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979145.

Bateman:1884:WPF

[207] A. E. Bateman. Wine production in France. Journal of the Statistical Society, 47(4):609–613, December 1884. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979146.

Jeans:1884:CEE

[208] J. S. Jeans. On the comparative efficiency and earnings of labour at home and abroad. Journal of the Statistical Society, 47(4):614–665, December 1884. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979147.

Anonymous:1884:ARY

[209] Anonymous. The agricultural returns for the year 1884. Journal of the Statistical Society, 47(4):666–691, December 1884. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979148.

Anonymous:1884:SBS

[210] Anonymous. Savings banks statistics. Journal of the Statistical Society, 47(4): 691–694, December 1884. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979149.

Anonymous:1884:IRC

[211] Anonymous. Industrial Remuneration Conference. Journal of the Statistical Society, 47(4):694, December 1884. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979150.

Anonymous:1884:FCS

[212] Anonymous. The forthcoming celebration of the Society’s jubilee. Journal of the Statistical Society, 47(4):695–696, December 1884. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979151.

Anonymous:1884:ALd

[213] Anonymous. Additions to the library.

Journal of the Statistical Society, 47(4): 696–712, December 1884. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979152.

Anonymous:1885:VIa

[214] Anonymous. Volume information. Journal of the Statistical Society, 48(1):i– viii, March 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979257.

Temple:1885:PSC

[215] Richard Temple. Population statistics of China. Journal of the Statistical Society, 48(1):1–20, March 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979258.

Brabrook:1885:RST

[216] E. W. Brabrook. The relation of the state to thrift; ten years’ statistics of friendly societies and similar institutions. Journal of the Statistical Society, 48(1):21–53, March 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979259.

Anonymous:1885:FCH

[217] Anonymous. Financial and commercial history of 1884. Journal of the Statistical Society, 48(1):54–63, March 1885. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979260.

Levasseur:1885:ENC

[218] Emile Levasseur. Emigration in the Nineteenth Century. Journal of the Statistical Society, 48(1):63–81, March 1885. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979261.

Roth:1885:ASC

[219] Henry Ling Roth. Arbere. a short contribution to the study of peasant proprietorship. Journal of the Statistical Society, 48(1):81–91, March 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979262.

Anonymous:1885:FLD

[220] Anonymous. The fires in London during the year 1884, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 48(1):91–96, March 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979263.

Anonymous:1885:GL

[221] Anonymous. German literature of 1883 and 1884. Journal of the Statistical Society, 48(1):97, March 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979264.

Anonymous:1885:EL

[222] Anonymous. English literature in 1884. Journal of the Statistical Society, 48(1): 97–99, March 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979265.

Anonymous:1885:EIY

[223] Anonymous. Emigration and immigration in the year 1884. Journal of the Statistical Society, 48(1):100–108, March 1885. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979266.

Dunn:1885:TM

[224] Archibald J. Dunn. The trade of Marseilles. Journal of the Statistical Society, 48(1):108–116, March 1885. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979267.

Anonymous:1885:SMU

[225] Anonymous. Statistics of murder in the United States. Journal of the Statistical Society, 48(1):116–121, March 1885. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979268.

Anonymous:1885:IRC

[226] Anonymous. The Industrial Remuneration Conference. Journal of the Statistical Society, 48(1):121, March 1885. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979269.

Anonymous:1885:ALa

[227] Anonymous. Additions to the library.

Journal of the Statistical Society, 48(1): 122–142, March 1885. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979270.

Anonymous:1885:PR

[228] Anonymous. Periodical returns. Journal of the Statistical Society, 48(1):143– 166, March 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979271.

Ravenstein:1885:LM

[229] E. G. Ravenstein. The laws of migration. Journal of the Statistical Society, 48(2):167–235, June 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979181.

Connell:1885:IRI

[230] Arthur Knatchbull Connell. Indian railways and Indian wheat. Journal of the Statistical Society, 48(2):236–276, June 1885. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979182.

Foster:1885:RDC

[231] C. Le Neve Foster. On the relative dangers of coal and metal mining in the United Kingdom of Great Britain and Ireland. Journal of the Statistical Society, 48(2):277–279, June 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979183.

Anonymous:1885:LSM

[232] Anonymous. Lloyd’s statistics of marine casualties for the year 1884. Journal of the Statistical Society, 48(2):280– 287, June 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979184.

Anonymous:1885:SRL

[233] Anonymous. Statistical results of the last census of France, Prussia, Austria, Italy, Switzerland, and Belgium. Journal of the Statistical Society, 48(2):288– 297, June 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979185.

Anonymous:1885:ACE

[234] Anonymous. The agricultural crisis in Europe. Journal of the Statistical Society, 48(2):297–312, June 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979186.

Anonymous:1885:ALb

[235] Anonymous. Additions to the library.

Journal of the Statistical Society, 48(2): 313–332, June 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979187.

Anonymous:1885:RCF

[236] Anonymous. Report of the Council for the financial year ended 31st December, 1884, and for the sessional year ending 23rd June, 1885, presented at the fiftyfirst annual meeting of the Statistical Society, held in the theatre of the University of London, Burlington Gardens, W., on the 23rd of June, 1885. Journal of the Statistical Society, 48(3):333–346, September 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979954.

Anonymous:1885:PFF

[237] Anonymous. Proceedings of the fiftyfirst anniversary meeting. Journal of the Statistical Society, 48(3):346–348, September 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979955.

Price-Williams:1885:PL

[238] R. Price-Williams. The population of London, 1801–81. Journal of the Statistical Society, 48(3):349–440, September 1885. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979956.

Anonymous:1885:RLE

[239] Anonymous. Recent legislation and economic science: The Newmarch Memorial Essay. Journal of the Statistical Society, 48(3):441–468, September 1885. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979957.

Beaulieu:1885:WPF

[240] Paul Leroy Beaulieu. Wine production in France. Journal of the Statistical Society, 48(3):469–482, September 1885. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979958.

Anonymous:1885:IRR

[241] Anonymous. Inland revenue report, 1884–85, with some retrospective history. Journal of the Statistical Society, 48(3):482–484, September 1885. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979959.

Anonymous:1885:FTU

[242] Anonymous. Foreign trade of the United Kingdom. Journal of the Statistical Society, 48(3):484–491, September 1885. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979960.

Anonymous:1885:SJ

[243] Anonymous. The Society’s jubilee. Journal of the Statistical Society, 48(3):491– 492, September 1885. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979961.

Anonymous:1885:ALc

[244] Anonymous. Additions to the library. Journal of the Statistical Society, 48(3): 492–504, September 1885. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979962.

Anonymous:1885:VIb

[245] Anonymous. Volume information. Journal of the Statistical Society, 48(4):699– 716, December 1885. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979196.

Rawson:1885:ISI

[246] Rawson W. Rawson. International statistics, illustrated by vital statistics of Europe and of some of the United States of America; being the opening address of Sir Rawson, K.C.M.G., C.B., President of the Statistical Society and of the International Statistical Institute. Session 1885–86. delivered 17th November, 1885. Journal of the Statistical Society, 48(4):505–592, December 1885. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979197.

Anonymous:1885:PN

[247] Anonymous. Proceedings on the 17th November, 1885. Journal of the Statistical Society, 48(4):593–594, December 1885. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979198.

Sidgwick:1885:ESS

[248] Henry Sidgwick. “Economic science and statistics.” The address of the President of Section F of the British Association, at the fifty-fifth meeting, held at Aberdeen, in September, 1885. Journal of the Statistical Society, 48(4):595–616, December 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979199.

Bateman:1885:CT

[249] A. E. Bateman. Customs tariffs. Journal of the Statistical Society, 48(4):617–627, December 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979200.

Edgeworth:1885:MAV

[250] F. Y. Edgeworth. On methods of ascertaining variations in the rate of births, deaths, and marriages. Journal of the Statistical Society, 48(4):628–649, December 1885. CODEN ???? ISSN

                                        0959-5341. URL http://www.jstor.
                                        org/stable/2979201.

Anonymous:1885:DWG

[251] Anonymous. Dr. William A. Guy. Journal of the Statistical Society, 48(4):650– 651, December 1885. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979202.

Anonymous:1885:MCW

[252] Anonymous. Mr. Cornelius Walford. Journal of the Statistical Society, 48(4): 651–652, December 1885. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979203.

Anonymous:1885:MPF

[253] Anonymous. The migrations of the population in France. Journal of the Statistical Society, 48(4):652–667, December 1885. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979204.

Anonymous:1885:SAC

[254] Anonymous. Statistics of Australasian colonies. Journal of the Statistical Society, 48(4):668–669, December 1885. CO-DEN ???? ISSN 0959-5341. URL http:

                                        //www.jstor.org/stable/2979205.

Anonymous:1885:SPG

[255] Anonymous. Small pox in Germany. Journal of the Statistical Society, 48(4): 670–674, December 1885. CODEN ????

ISSN 0959-5341. URL http://www.

jstor.org/stable/2979206.

Anonymous:1885:RGT

[256] Anonymous. The Registrar-General’s ten years’ supplement. Journal of the Statistical Society, 48(4):674–680, December 1885. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979207.

Anonymous:1885:ALd

[257] Anonymous. Additions to the library.

Journal of the Statistical Society, 48(4): 681–698, December 1885. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979208.

Anonymous:1886:VIa

[258] Anonymous. Volume information. Journal of the Statistical Society, 49(1):i– viii, March 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979224.

Scrutton:1886:PLL

[259] Thomas Scrutton. “Preventible loss of life at sea”. Journal of the Statistical Society, 49(1):1–17 + 19–27, March 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979225.

Anonymous:1886:PD

[260] Anonymous. Proceedings on the 15th December, 1885. Journal of the Statistical Society, 49(1):17–18, March 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979226.

Giffen:1886:FNP

[261] Robert Giffen. Further notes on the progress of the working classes in the last half century. Journal of the Statistical Society, 49(1):28–100, March 1886. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979227.

Ogle:1886:SEW

[262] William Ogle. Suicides in England and Wales in relation to age, sex, season, and occupation. Journal of the Statistical Society, 49(1):101–135, March 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979228.

Anonymous:1886:FCH

[263] Anonymous. Financial and commercial history of 1885. Journal of the Statistical Society, 49(1):136–144, March 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979229.

Anonymous:1886:ARY

[264] Anonymous. The agricultural returns for the year 1885. Journal of the Statistical Society, 49(1):144–163, March 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979230.

Anonymous:1886:MMP

[265] Anonymous. Mortality in the medical profession. Journal of the Statistical Society, 49(1):164–170, March 1886. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979231.

Anonymous:1886:FLD

[266] Anonymous. The fires in London during the year 1885, and the Metropolitan Fire Brigade. Journal of the Statistical Society, 49(1):170–175, March 1886. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979232.

Anonymous:1886:GL

[267] Anonymous. German literature of 1884 and 1885. Journal of the Statistical Society, 49(1):176, March 1886. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979233.

Anonymous:1886:EL

[268] Anonymous. English literature in 1885.

Journal of the Statistical Society, 49(1): 176–178, March 1886. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979234.

Anonymous:1886:PPM

[269] Anonymous. The production of precious metals and the monetary circulation in various countries. Journal of the Statistical Society, 49(1):178–186, March 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979235.

deFlaix:1886:NWF

[270] E. Fournier de Flaix. The national wealth of France compared with other countries. Journal of the Statistical Society, 49(1):186–200, March 1886. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979236.

Anonymous:1886:ALa

[271] Anonymous. Additions to the library.

Journal of the Statistical Society, 49(1): 201–216, March 1886. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979237.

Anonymous:1886:PR

[272] Anonymous. Periodical returns. Journal of the Statistical Society, 49(1):217– 240, March 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979238.

Levi:1886:PJS

[273] Leone Levi. “The progress of joint stock companies with limited and unlimited liability in the United Kingdom, during the “fifteen years 1869–84.” (in continuation of a paper read before the Society in January, 1870). Journal of the Statistical Society, 49(2):241–272, June 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979153.

Bourne:1886:FCT

[274] Stephen Bourne. Foreign and colonial tariffs as influencing prices and affecting trade. Journal of the Statistical Society, 49(2):273–313, June 1886. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979154.

Booth:1886:OPU

[275] Charles Booth. Occupations of the people of the United Kingdom, 1801–81. Journal of the Statistical Society, 49(2): 314–444, June 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979155.

Anonymous:1886:EIY

[276] Anonymous. Emigration and immigration in the year 1885. Journal of the Statistical Society, 49(2):445–450, June 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979156.

deBeaucaire:1886:MRD

[277] M. Le Vicomte Rorric de Beaucaire. The migration from the rural districts and the condition of the agricultural population in Germany. Journal of the Statistical Society, 49(2):450–468, June 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979157.

Edgeworth:1886:PM

[278] F. Y. Edgeworth. Progressive means.

Journal of the Statistical Society, 49(2): 469–475, June 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979158.

Anonymous:1886:CPN

[279] Anonymous. Census of part of the North-West territories of Canada, August, 1885. Journal of the Statistical Society, 49(2):476–479, June 1886. CO-DEN ???? ISSN 0959-5341. URL http:

//www.jstor.org/stable/2979159.

Anonymous:1886:ICL

[280] Anonymous. Index to the catalogue of the library. Journal of the Statistical Society, 49(2):480, June 1886. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979160.

Anonymous:1886:SNP

[281] Anonymous. Statistics of the newspaper and periodical Press. Journal of the Statistical Society, 49(2):480–482, June 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979161.

Anonymous:1886:NSR

[282] Anonymous. Notes on some recent additions to the library. Journal of the Statistical Society, 49(2):482–494, June 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979162.

Anonymous:1886:ALb

[283] Anonymous. Additions to the library. Journal of the Statistical Society, 49(2): 494–512, June 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979163.

Rawson:1886:RCF

[284] Rawson W. Rawson. Report of the Council for the financial year ended 31st December, 1885, and for the sessional year ending 29th June, 1886, presented at the fifty-second annual meeting of the Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 29th of June, 1886. Journal of the Statistical Society, 49(3):513–531, September 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979291.

Anonymous:1886:PFS

[285] Anonymous. Proceedings of the fiftysecond anniversary meeting. Journal of the Statistical Society, 49(3):532–538, September 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979292.

Stout:1886:NPN

[286] Robert Stout. Notes on the progress of New Zealand for twenty years, 1864– 84. Journal of the Statistical Society, 49 (3):539–580, September 1886. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979293.

Sauerbeck:1886:PCP

[287] Augustus Sauerbeck. Prices of commodities and the precious metals. Journal of the Statistical Society, 49(3):581– 648, September 1886. CODEN ????

ISSN 0959-5341. URL http://www.

jstor.org/stable/2979294.

Edgeworth:1886:MMS

[288] F. Y. Edgeworth. The mathematical method of statistics. Journal of the Statistical Society, 49(3):649–654, September 1886. CODEN ???? ISSN 0959-

5341. URL http://www.jstor.org/
stable/2979295.

Anonymous:1886:CIC

[289] Anonymous. On a calculating instrument called a universal proportion table. Journal of the Statistical Society, 49 (3):655–656, September 1886. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979296.

Anonymous:1886:LSM

[290] Anonymous. Lloyd’s statistics of marine casualties for 1885. Journal of the Statistical Society, 49(3):657–661, September 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979297.

Loua:1886:SRP

[291] M. Toussaint Loua. Statistical review of the present position of Italy. Journal of the Statistical Society, 49(3):662–672, September 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979298.

Anonymous:1886:NESa

[292] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 49(3):672–678, September 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979299.

Anonymous:1886:ALc

[293] Anonymous. Additions to the library. Journal of the Statistical Society, 49(3): 678–692, September 1886. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979300.

Anonymous:1886:VIb

[294] Anonymous. Volume information. Journal of the Statistical Society, 49(4):805– 823, December 1886. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979164.

Jeans:1886:CCW

[295] J. S. Jeans. On the cost and the conditions of working railway traffic in different countries. Journal of the Statistical Society, 49(4):693–735, December 1886. CODEN ???? ISSN 0959-5341.

                                        URL http://www.jstor.org/stable/
                                        2979165.

Martin:1886:ESS

[296] John Biddulph Martin. “Economic science and statistics.” The address of the President of Section F of the British Association, at the fifty-sixth meeting, held at Birmingham, in September, 1886. Journal of the Statistical Society, 49(4): 736–754, December 1886. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979166.

Acland:1886:WMC

[297] Arthur H. D. Acland. Working men’s cooperative organizations in Great Britain. Journal of the Statistical Society, 49(4): 755–759, December 1886. CODEN ???? ISSN 0959-5341. URL http://www.

                                        jstor.org/stable/2979167.

Longstaff:1886:PPU

[298] G. B. Longstaff. The present population of the United Kingdom, 1st April, 1886. Journal of the Statistical Society, 49(4):760–763, December 1886. CODEN ???? ISSN 0959-5341. URL http://

www.jstor.org/stable/2979168.

DeFlaix:1886:ISD

[299] E. Fournier De Flaix. The increase of shipping and Dock accomodation in the principal maritime ports. Journal of the Statistical Society, 49(4):763–776, December 1886. CODEN ???? ISSN

0959-5341. URL http://www.jstor.
org/stable/2979169.

Anonymous:1886:NESb

[300] Anonymous. Notes on economical and statistical works. Journal of the Statistical Society, 49(4):776–783, December 1886. CODEN ???? ISSN 0959-5341.

URL http://www.jstor.org/stable/
2979170.

Anonymous:1886:ALd

[301] Anonymous. Additions to the library.

Journal of the Statistical Society, 49(4): 783–804, December 1886. CODEN ???? ISSN 0959-5341. URL http://www.

jstor.org/stable/2979171.

Anonymous:1887:VIa

[302] Anonymous. Volume information. Journal of the Royal Statistical Society, 50 (1):i–viii, March 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979118.

Anonymous:1887:FM

[303] Anonymous. Front matter. Journal of the Royal Statistical Society, 50(1): ii–iii, March 1887. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979119.

Anonymous:1887:CRS

[304] Anonymous. [charter of] the Royal Statistical Society. Journal of the Royal Statistical Society, 50(1):1–4, March 1887. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979120.

Price:1887:SSO

[305] L. L. F. R. Price. “Sliding scales and other methods of wage-arrangement in the North of England”. Journal of the Royal Statistical Society, 50(1): 5–85, March 1887. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979121.

Craigie:1887:SDA

[306] P. G. Craigie. The size and distribution of agricultural holdings in England and abroad. Journal of the Royal Statistical Society, 50(1):86–149, March 1887. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979122.

Nicholson:1887:MVV

[307] J. Shield Nicholson. The measurement of variations in the value of the monetary standard. Journal of the Royal Statistical Society, 50(1):150–166, March 1887. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979123.

Anonymous:1887:RSS

[308] Anonymous. The Royal Statistical Society. Journal of the Royal Statistical Society, 50(1):167, March 1887. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979124.

Anonymous:1887:II

[309] Anonymous. The Imperial Institute.

Journal of the Royal Statistical Society, 50(1):167–169, March 1887. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979125.

Anonymous:1887:SICa

[310] Anonymous. The Sixth International Congress of Hygiene and Demography. Journal of the Royal Statistical Society, 50(1):169, March 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979126.

Anonymous:1887:FCH

[311] Anonymous. Financial and commercial history of 1886. Journal of the Royal Statistical Society, 50(1):169– 180, March 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979127.

Anonymous:1887:PCa

[312] Anonymous. Prices of commodities.

Journal of the Royal Statistical Society, 50(1):180–182, March 1887. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979128.

Anonymous:1887:DFP

[313] Anonymous. The decline of the French population (a translation). Journal of the Royal Statistical Society, 50(1):183– 197, March 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979129.

Anonymous:1887:FLD

[314] Anonymous. The fires in London during the year 1886, and the Metropolitan Fire Brigade. Journal of the Royal Statistical Society, 50(1):197–201, March 1887. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979130.

Anonymous:1887:EL

[315] Anonymous. English literature in 1886. Journal of the Royal Statistical Society, 50(1):202–203, March 1887. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979131.

Anonymous:1887:ISI

[316] Anonymous. International Statistical Institute. Journal of the Royal Statistical Society, 50(1):204–205, March 1887. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979132.

Anonymous:1887:NESa

[317] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 50(1):205– 217, March 1887. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979133.

Anonymous:1887:ALa

[318] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 50(1):217–230, March 1887. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979134.

Anonymous:1887:PR

[319] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 50(1): 231–254, March 1887. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979135.

Humphreys:1887:CMS

[320] Noel A. Humphreys. “Class mortality statistics. Journal of the Royal Statistical Society, 50(2):255–292, June 1887.

ISSN 0952-8385. URL http://www.

jstor.org/stable/2340189.

Elliott:1887:ATP

[321] Thomas Henry Elliott. The annual taxes on property and income. Journal of the Royal Statistical Society, 50(2):293–325, June 1887. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340190.

Booth:1887:ITH

[322] Charles Booth. The inhabitants of tower hamlets (school board division), their condition and occupations. Journal of the Royal Statistical Society, 50(2): 326–401, June 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2340191.

Anonymous:1887:EIY

[323] Anonymous. Emigration and immigration in the year 1886. Journal of the Royal Statistical Society, 50(2):402–408, June 1887. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340192.

Anonymous:1887:EID

[324] Anonymous. Effects of import duties in new and old countries. Journal of the Royal Statistical Society, 50(2):408–410, June 1887. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2340193.

Anonymous:1887:NESb

[325] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 50(2):410–420, June 1887. ISSN 0952-8385. URL http://

www.jstor.org/stable/2340194.

Anonymous:1887:ALb

[326] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 50(2):420–436, June 1887. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2340195.

Balfour:1887:RCF

[327] T. Graham Balfour. Report of the Council for the financial year ended 31st December, 1886, and for the sessional year ending 28th June, 1887, presented at the fifty-third annual meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 28th of June, 1887. Journal of the Royal Statistical Society, 50(3): 437–449, September 1887. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979059.

Anonymous:1887:PFT

[328] Anonymous. Proceedings of the fiftythird anniversary meeting. Journal of the Royal Statistical Society, 50(3):450– 457, September 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979060.

Lawson:1887:RMA

[329] Robert Lawson. Remarks on the mortality among the troops serving in the United Kingdom from consumption. Journal of the Royal Statistical Society, 50(3):458–494, September 1887. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979061.

Rabino:1887:SSS

[330] Joseph Rabino. The statistical story of the Suez Canal. Journal of the Royal Statistical Society, 50(3):495–546, September 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979062.

Levasseur:1887:TMS

[331] Emile Levasseur. The tables of mortality and survivorship. Journal of the Royal Statistical Society, 50(3):547– 569, September 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979063.

Anonymous:1887:OSJ

[332] Anonymous. The official statistics of Japan. Journal of the Royal Statistical Society, 50(3):569–571, September 1887. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979064.

Anonymous:1887:NESc

[333] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 50(3):571–576, September 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979065.

Anonymous:1887:ALc

[334] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 50(3):576–588, September 1887. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979066.

Anonymous:1887:VIb

[335] Anonymous. Volume information. Journal of the Royal Statistical Society, 50 (4):769–789, December 1887. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979373.

Goschen:1887:IMI

[336] G. J. Goschen. “The increase of moderate incomes;” being the inaugural address of the President of the Royal Statistical Society, the Right Hon. G. J.

Goschen, M.P., Chancellor of the Exchequer. Session 1887–88. delivered 6th December, 1887 at Willis’s rooms. Journal of the Royal Statistical Society, 50 (4):589–612, December 1887. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979374.

Anonymous:1887:PD

[337] Anonymous. Proceedings on the 6th December, 1887. Journal of the Royal Statistical Society, 50(4):613–614, December 1887. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979375.

Giffen:1887:RRM

[338] Robert Giffen. “The recent rate of material progress in England”: The address of the President of Section F. — Economic Science and Statistics — of the British Association, at the fifty-seventh meeting, held at Manchester, September, 1887. Journal of the Royal Statistical Society, 50(4):615–647, December 1887. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979376.

Ogle:1887:SSM

[339] William Ogle. Summary of several male life tables. Journal of the Royal Statistical Society, 50(4):648–652, December 1887. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979377.

Bateman:1887:OSF

[340] A. E. Bateman. Our statistics of foreign trade, and what they tell us. Journal of the Royal Statistical Society, 50(4):653– 658, December 1887. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979378.

Levi:1887:PCI

[341] Leone Levi. On the progress of commerce and industry during the last fifty years. Journal of the Royal Statistical Society, 50(4):659–668, December 1887. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979379.

Bernard:1887:WWP

[342] Francois Bernard. The world’s wheat production. Journal of the Royal Statistical Society, 50(4):668–688, December 1887. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979380.

Anonymous:1887:PCb

[343] Anonymous. The population of China.

Journal of the Royal Statistical Society, 50(4):688–696, December 1887. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979381.

Anonymous:1887:SICb

[344] Anonymous. The sixth International Congress of Hygiene and Demography. Journal of the Royal Statistical Society, 50(4):696–703, December 1887. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979382.

Anonymous:1887:AS

[345] Anonymous. The abuse of statistics.

Journal of the Royal Statistical Society, 50(4):703–708, December 1887. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979383.

Anonymous:1887:AR

[346] Anonymous. Agricultural returns for 1887. Journal of the Royal Statistical Society, 50(4):708–725, December 1887. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979384.

Anonymous:1887:CDF

[347] Anonymous. Cattle and dairy farming of the world. Journal of the Royal Statistical Society, 50(4):725–726, December 1887. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979385.

Anonymous:1887:TI

[348] Anonymous. The trade improvements in 1887. Journal of the Royal Statistical Society, 50(4):726–727, December 1887. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979386.

Anonymous:1887:NESd

[349] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 50(4):727– 744, December 1887. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979387.

Anonymous:1887:ALd

[350] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 50(4):745–768, December 1887. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979388.

Anonymous:1888:VIa

[351] Anonymous. Volume information. Journal of the Royal Statistical Society, 51 (1):i–viii, March 1888. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979078.

Anonymous:1888:EQS

[352] Anonymous. Errata: Quantitation of stock exchange values. Journal of the Royal Statistical Society, 51(1): iv, March 1888. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979079.

Anonymous:1888:E

[353] Anonymous. Errata. Journal of the Royal Statistical Society, 51(1): iv, March 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979080.

Danvers:1888:DER

[354] Juland Danvers. Defects of English railway statistics. Journal of the Royal Statistical Society, 51(1):1–32, March 1888. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979081.

Jones:1888:POA

[355] Benjamin Jones. Progress, organisation, and aims of working class co-operators. Journal of the Royal Statistical Society, 51(1):33–73, March 1888. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979082.

Colmer:1888:SCR

[356] Joseph G. Colmer. Some Canadian railway and commercial statistics. Journal of the Royal Statistical Society, 51(1): 74–112, March 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979083.

Edgeworth:1888:MTB

[357] F. Y. Edgeworth. The mathematical theory of banking. Journal of the Royal Statistical Society, 51(1):113– 127, March 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979084.

Anonymous:1888:FCH

[358] Anonymous. Financial and commercial history of 1887. Journal of the Royal Statistical Society, 51(1):127– 136, March 1888. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979085.

Anonymous:1888:FLD

[359] Anonymous. Fires in London during the year 1887, and the Metropolitan Fire Brigade. Journal of the Royal Statistical Society, 51(1):136–140, March 1888. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979086.

Raffalovich:1888:ESB

[360] M. Arthur Raffalovich. The effects of shipping bounties. Journal of the Royal Statistical Society, 51(1):141– 154, March 1888. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979087.

Sauerbeck:1888:PC

[361] Augustus Sauerbeck. Prices of commodities in 1886 and 1887. Journal of the Royal Statistical Society, 51(1): 154–157, March 1888. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979088.

Anonymous:1888:DIG

[362] Anonymous. Distribution of incomes in Germany. Journal of the Royal Statistical Society, 51(1):158–161, March 1888. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979089.

Anonymous:1888:USO

[363] Anonymous. Unsuitable subjects for official statistics. Journal of the Royal Statistical Society, 51(1):161–162, March 1888. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979090.

Anonymous:1888:AS

[364] Anonymous. The abuse of statistics. Journal of the Royal Statistical Society, 51(1):162–163, March 1888. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979091.

Anonymous:1888:GL

[365] Anonymous. German literature in 1886 and 1887. Journal of the Royal Statistical Society, 51(1):163–164, March 1888. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979092.

Anonymous:1888:EL

[366] Anonymous. English literature in 1887.

Journal of the Royal Statistical Society, 51(1):164–166, March 1888. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979093.

Anonymous:1888:ALa

[367] Anonymous. American literature in 1886 and 1887. Journal of the Royal Statistical Society, 51(1):166, March 1888. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979094.

Anonymous:1888:WBP

[368] Anonymous. The world’s big ports.

Journal of the Royal Statistical Society, 51(1):167–169, March 1888. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979095.

Anonymous:1888:NESa

[369] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 51(1):169– 186, March 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979096.

Anonymous:1888:ALb

[370] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 51(1):186–200, March 1888. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979097.

Anonymous:1888:PR

[371] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 51(1): 201–224, March 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979098.

Palgrave:1888:PTL

[372] R. H. Inglis Palgrave. Progressive taxation as levied in Switzerland. Cantons Basel-Stadt, de Vaud, and Uri. Journal of the Royal Statistical Society, 51(2): 225–275, June 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979108.

Booth:1888:COP

[373] Charles Booth. Condition and occupations of the people of East London and Hackney, 1887. Journal of the Royal Statistical Society, 51(2):276–339, June 1888. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979109.

Levi:1888:PLL

[374] Leone Levi. Professor Leone Levi, LL.D. Journal of the Royal Statistical Society, 51(2):340–342, June 1888. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979110.

vonNeumann-Spallart:1888:DFX

[375] F. X. von Neumann-Spallart. Dr. F. X. von Neumann-Spallart. Journal of the Royal Statistical Society, 51(2):343–346, June 1888. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979111.

Edgeworth:1888:SNM

[376] F. Y. Edgeworth. Some new methods of measuring variation in general prices. Journal of the Royal Statistical Society, 51(2):346–368, June 1888. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979112.

Anonymous:1888:EIY

[377] Anonymous. Emigration and immigration in the year 1887. Journal of the Royal Statistical Society, 51(2):368–375, June 1888. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979113.

Anonymous:1888:RAR

[378] Anonymous. Relative advantages of the railways and waterways of Germany. Journal of the Royal Statistical Society, 51(2):375–396, June 1888. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979114.

Anonymous:1888:SRE

[379] Anonymous. Statistics of the Russian empire. Journal of the Royal Statistical Society, 51(2):396–398, June 1888. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979115.

Anonymous:1888:NESb

[380] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 51(2):398–414, June 1888. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979116.

Anonymous:1888:ALc

[381] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 51(2):414–432, June 1888. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979117.

Kennedy:1888:RCF

[382] C. M. Kennedy. Report of the Council for the financial year ended 31st December, 1887, and for the sessional year ending 26th June, 1888, presented at the fifty-fourth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 26th of June, 1888. Journal of the Royal Statistical Society, 51(3):433–447, September 1888. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339893.

Anonymous:1888:PFF

[383] Anonymous. Proceedings of the fiftyfourth Annual General Meeting. Journal of the Royal Statistical Society, 51(3): 448–449, September 1888. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2339894.

Craigie:1888:EPR

[384] P. G. Craigie. The English poor rate: Some recent statistics of its administration and pressure. Journal of the Royal Statistical Society, 51(3):450– 493, September 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339895.

Price:1888:WBN

[385] L. L. Price. “West Barbary;” or, notes on the system of work and wages in the Cornish Mines. Journal of the Royal Statistical Society, 51(3):494–566, September 1888. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2339896.

Ellis:1888:QSE

[386] Arthur Ellis. The “quantitation” of stock exchange values. Journal of the Royal Statistical Society, 51(3):567– 598, September 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339897.

Edgeworth:1888:SE

[387] F. Y. Edgeworth. The statistics of examinations. Journal of the Royal Statistical Society, 51(3):599–635, September 1888. ISSN 0952-8385. URL http://

www.jstor.org/stable/2339898.

Anonymous:1888:SIP

[388] Anonymous. Statistics of incomes in Prussia, Saxony, and the United Kingdom. Journal of the Royal Statistical Society, 51(3):635–649, September 1888. ISSN 0952-8385. URL http://www.

jstor.org/stable/2339899.

Anonymous:1888:APS

[389] Anonymous. Agriculture and prices in the Seventeenth Century. Journal of the Royal Statistical Society, 51(3):649– 654, September 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339900.

Anonymous:1888:NESc

[390] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 51(3):654–664, September 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2339901.

Anonymous:1888:ALd

[391] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 51(3):664–682, September 1888. ISSN

0952-8385. URL http://www.jstor.
org/stable/2339902.

Anonymous:1888:VIb

[392] Anonymous. Volume information. Journal of the Royal Statistical Society, 51 (4):855–877, December 1888. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979067.

Balfour:1888:IAD

[393] T. Graham Balfour. The inaugural address of Dr. T. Graham Balfour, F.R.S., &c., honorary physician to Her Majesty the Queen, President of the Royal Statistical Society. Delivered 20th November, 1888. Journal of the Royal Statistical Society, 51(4):683–700, December 1888. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979068.

Anonymous:1888:PN

[394] Anonymous. Proceedings on the 20th November, 1888. Journal of the Royal Statistical Society, 51(4):701– 703, December 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979069.

Bramwell:1888:APS

[395] Lord Bramwell. The address of the President of Section F — Economic Science and Statistics — of the British Association, at the fifty-eighth meeting, held at Bath, in September, 1888. Journal of the Royal Statistical Society, 51(4):704– 712, December 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979070.

Giffen:1888:RCP

[396] R. Giffen. Recent changes in prices and incomes compared. Journal of the Royal Statistical Society, 51(4):713– 815, December 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979071.

Anonymous:1888:RCR

[397] Anonymous. Report of a committee with reference to the census of 1891. Journal of the Royal Statistical Society, 51 (4):816–818, December 1888. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979072.

Bourne:1888:INI

[398] Stephen Bourne. Index numbers as illustrating the progressive exports of British produce and manufactures. Journal of the Royal Statistical Society, 51(4):819– 823, December 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979073.

Anonymous:1888:WWC

[399] Anonymous. The wages of the working classes in Berlin. Journal of the Royal Statistical Society, 51(4):823– 826, December 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979074.

Anonymous:1888:AMD

[400] Anonymous. Awards for meritorious discoveries and inventions. Journal of the Royal Statistical Society, 51(4):826– 827, December 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979075.

Anonymous:1888:NESd

[401] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 51(4):827– 838, December 1888. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979076.

Anonymous:1888:ALe

[402] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 51(4):838–854, December 1888. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979077.

Anonymous:1889:VIa

[403] Anonymous. Volume information. Journal of the Royal Statistical Society, 52 (1):i–viii, March 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979361.

Price-Williams:1889:CQ

[404] Richard Price-Williams. The coal question. Journal of the Royal Statistical Society, 52(1):1–46, March 1889. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979362.

Palgrave:1889:LGD

[405] R. H. Inglis Palgrave. Local government. On the distribution of the licenses proposed to be transferred in aid of local expenditure. Journal of the Royal Statistical Society, 52(1):47–59, March 1889. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979363.

Sorley:1889:MRT

[406] W. R. Sorley. Mining royalties and their effect on the iron and coal trades. Journal of the Royal Statistical Society, 52 (1):60–98, March 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979364.

deHaas:1889:TER

[407] Jacob de Haas. A third element in the rate of interest. Journal of the Royal Statistical Society, 52(1):99–116, March 1889. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979365.

Sauerbeck:1889:PC

[408] A. Sauerbeck. Prices of commodities in 1887 and 1888. Journal of the Royal Statistical Society, 52(1):116–120, March 1889. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979366.

Anonymous:1889:DDV

[409] Anonymous. Death duties in various countries. Journal of the Royal Statistical Society, 52(1):120–143, March 1889. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979367.

Anonymous:1889:FLD

[410] Anonymous. Fires in London during the year 1888, and the Metropolitan Fire Brigade. Journal of the Royal Statistical Society, 52(1):144–148, March 1889. ISSN 0952-8385. URL http://www.

jstor.org/stable/2979368.

Anonymous:1889:EL

[411] Anonymous. English literature in 1888.

Journal of the Royal Statistical Society, 52(1):148–150, March 1889. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979369.

Anonymous:1889:NESa

[412] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 52(1):150– 163, March 1889. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979370.

Anonymous:1889:ALa

[413] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 52(1):164–180, March 1889. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979371.

Anonymous:1889:PR

[414] Anonymous. Periodical returns. Journal of the Royal Statistical Society, 52(1): 181–204, March 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979372.

Ogle:1889:ADR

[415] William Ogle. The alleged depopulation of the rural districts of England. Journal of the Royal Statistical Society, 52(2): 205–240, June 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979332.

Ravenstein:1889:LM

[416] E. G. Ravenstein. The laws of migration. Journal of the Royal Statistical Society, 52(2):241–305, June 1889. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979333.

Neymarck:1889:FEC

[417] Alfred Neymarck. France, 1789–1889. an economic centenary. Journal of the Royal Statistical Society, 52(2):306–338, June 1889. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979334.

Anonymous:1889:GL

[418] Anonymous. German literature in 1887 and 1888. Journal of the Royal Statistical Society, 52(2):338–339, June 1889. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979335.

Anonymous:1889:EIY

[419] Anonymous. Emigration and immigration in the year 1888. Journal of the Royal Statistical Society, 52(2):339–345, June 1889. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979336.

Anonymous:1889:CQB

[420] Anonymous. The coal question:

“Bunker” coal consumed. Journal of the Royal Statistical Society, 52(2):345–349, June 1889. ISSN 0952-8385. URL http:

//www.jstor.org/stable/2979337.

Anonymous:1889:NESb

[421] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 52(2):349–357, June 1889. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979338.

Anonymous:1889:ALb

[422] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 52(2):357–376, June 1889. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979339.

Anonymous:1889:RCF

[423] Anonymous. Report of the Council for the financial year ended 31st December, 1888, and for the sessional year ending 27th June, 1889, presented at the fifty-fifth Annual General Meeting of the Royal Statistical Society, held at the Society’s rooms, 9, Adelphi Terrace, Strand, London, on the 27th of June, 1889. Journal of the Royal Statistical Society, 52(3):377–395, September 1889. ISSN 0952-8385. URL http://

www.jstor.org/stable/2979136.

Anonymous:1889:PFF

[424] Anonymous. Proceedings of the fiftyfifth Annual General Meeting. Journal of the Royal Statistical Society, 52(3):

396–398, September 1889. ISSN 0952-

                                        8385. URL http://www.jstor.org/
                                        stable/2979137.

Bourne:1889:VVV

[425] Stephen Bourne. On variations in the volume and value of exports and imports of the United Kingdom in recent years. Journal of the Royal Statistical Society, 52(3):399–435, September 1889. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979138.

Longstaff:1889:SC

[426] G. B. Longstaff. Suggestions for the census of 1891. Journal of the Royal Statistical Society, 52(3):436–467, September 1889. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979139.

Anonymous:1889:SEC

[427] Anonymous. Sir Edwin Chadwick on an annual census. Journal of the Royal Statistical Society, 52(3):468–474, September 1889. ISSN 0952-8385. URL http:

                                        //www.jstor.org/stable/2979140.

Neymarck:1889:VPP

[428] M. A. Neymarck. The value of personal property in France. Journal of the Royal Statistical Society, 52(3):475– 493, September 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979141.

Anonymous:1889:NESc

[429] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 52(3):493–500, September 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979142.

Anonymous:1889:ALc

[430] Anonymous. Additions to the library.

Journal of the Royal Statistical Society, 52(3):501–516, September 1889. ISSN

0952-8385. URL http://www.jstor.
org/stable/2979143.

Anonymous:1889:VIb

[431] Anonymous. Volume information. Journal of the Royal Statistical Society, 52 (4):631–642, December 1889. ISSN 0952-

8385. URL http://www.jstor.org/
stable/2979099.

Balfour:1889:OAD

[432] T. Graham Balfour. The opening address of Dr. T. Graham Balfour, F.R.S., &c, Honorary Physician to Her Majesty the Queen, President of the Royal Statistical Society. Session 1889–90. delivered 19th November, 1889. Journal of the Royal Statistical Society, 52(4):517– 534, December 1889. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979100.

Anonymous:1889:PN

[433] Anonymous. Proceedings on the 19th November, 1889. Journal of the Royal Statistical Society, 52(4):535– 537, December 1889. ISSN 0952-8385.

URL http://www.jstor.org/stable/
2979101.

Edgeworth:1889:AMP

[434] F. Y. Edgeworth. “On the Application of Mathematics to Political Economy.” the address of the President of Section F — Economic Science and Statistics — of the British Association, at the fiftyninth meeting, held at Newcastle-Upon-Tyne, in September, 1889. Journal of the Royal Statistical Society, 52(4):538– 576, December 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979102.

Coste:1889:IMC

[435] Adolphe Coste. The International Monetary Congress of 1889. Journal of the Royal Statistical Society, 52(4):577– 595, December 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979103.

Anonymous:1889:PSV

[436] Anonymous. Public sympathy v. political economy. Journal of the Royal Statistical Society, 52(4):595–604, December 1889. ISSN 0952-8385. URL http://

                                        www.jstor.org/stable/2979104.

Anonymous:1889:DJS

[437] Anonymous. Directors of joint stock companies. Journal of the Royal Statistical Society, 52(4):604, December 1889. ISSN 0952-8385. URL http://www.

                                        jstor.org/stable/2979105.

Anonymous:1889:NESd

[438] Anonymous. Notes on economical and statistical works. Journal of the Royal Statistical Society, 52(4):605– 611, December 1889. ISSN 0952-8385.

                                        URL http://www.jstor.org/stable/
                                        2979106.

Anonymous:1889:ALd

[439] Anonymous. Additions to the library. Journal of the Royal Statistical Society, 52(4):611–630, December 1889. ISSN

                                        0952-8385. URL http://www.jstor.
                                        org/stable/2979107.